### v3

# A Non-Relativistic Limit for Heterotic Supergravity and its Gauge Lagrangian

Abstract
We show that the $D=10$ heterotic supergravity in a non-relativistic (NR) limit has a finite Lagrangian due to non-trivial cancellations of divergent parts arising from the Chern-Simons terms in the curvature of the $\hat{B}$-field and the Yang-Mills Lagrangian. This is similar to what happens in bosonic supergravity between the Ricci scalar, $\hat{R}$, and the $-\frac{1}{12}\hat{H}^{2}$ term, under the same limit. In this work we present the explicit form of the gauge transformations and curvatures after considering a NR limit and we compute the full finite Lagrangian in its covariant form. Particularly, we address both the ungauged part and the gauged part, and we write it using covariant derivatives and curvatures for the gauge fields. These results constitutes the first computation of a finite Lagrangian for heterotic supergravity under these kinds of limits. As an interesting feature of the expansion implemented, the Green-Schwarz mechanism for the two-form can be trivialized and terms equivalent to Chern-Simons contributions naturally arise from the previous property.
1 Introduction
In the last decade there has been huge progress made in the study of the different formulations of non-relativistic string theory [1]-[3] and its supergravity limit [4]-[17]  [^1] See [18]-[20] for reviews on this topic.. The low energy limit of bosonic string theory is given by the universal NS-NS supergravity, whose fundamental fields $\hat{g}$, $\hat{B}$, $\hat{\phi}$ admit a non-relativistic expansion of the form [13],
$\hat{g}_{\mu\nu}$ $=$ $c^{2}{\tau_{\mu}}^{a}{\tau_{\nu}}^{b}\eta_{ab}+h_{\mu\nu}\,,$ (1.1)
$\hat{g}^{\mu\nu}$ $=$ $\frac{1}{c^{2}}{\tau^{\mu}}_{a}{\tau^{\nu}}_{b}\eta^{ab}+h^{\mu\nu}\,,$ (1.2)
$\hat{B}_{\mu\nu}$ $=$ $-c^{2}{\tau_{\mu}}^{a}{\tau_{\nu}}^{b}\epsilon_{ab}+b_{\mu\nu}\,,$ (1.3)
$\hat{\phi}$ $=$ $\ln(c)+\varphi\,,$ (1.4)
where $\mu,\nu=1,\dots,10$ are space-time indices and we have split the flat index $\hat{a}=(a,a^{\prime})$ where $a,b,...=0,1$ are the transverse directions and $a^{\prime},b^{\prime},...=2,\dots,9$. The quantities in the previous expansion obey the following Newton-Cartan relations,
$\tau_{\mu}{}^{a}e^{\mu}{}_{a^{\prime}}$ $=\tau^{\mu}{}_{a}e_{\mu}{}^{a^{\prime}}=0\,,\qquad e_{\mu}{}^{a^{\prime}}e^{\mu}{}_{b^{\prime}}=\delta^{a^{\prime}}_{b^{\prime}}\,,$ (1.5)
$\tau_{\mu}{}^{a}\tau^{\mu b}$ $=\eta^{ab}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (1.6)
While, apriori, this expansion produces divergent tensorial quantities, i.e., Riemann tensor and its traces, H strength, etc., the NS-NS supergravity Lagrangian,
$S=\int d^{10}x\sqrt{-\hat{g}}e^{-2\hat{\phi}}\Big{(}\hat{R}+4\partial_{\mu}\hat{\phi}\partial^{\mu}\hat{\phi}-\frac{1}{12}\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho}\Big{)}\,,$ (1.7)
with
$\hat{H}_{\mu\nu\rho}=3\partial_{[\mu}\hat{B}_{\nu\rho]}\,,$ (1.8)
remains finite due to a curious cancellation between the divergent part of the Ricci scalar and the divergent part of the $\hat{H}^{2}$ term [13].
In the case of the bosonic sector of heterotic supergravity, we expect a similar cancellation when including the gauge field $\hat{A}_{\mu}^{i}$, $i=1,\dots,N$ and $N=496$ which is the dimension of the heterotic gauge group. Given a particular expansion for the gauge field, the inclusion of Chern-Simons terms in the curvature of the $\hat{B}$-field should produce divergences that must be compensated by the Yang-Mills Lagrangian. While there are several candidates for the expansion of the A-field in the heterotic scenario [21]-[22], there is still no explicit computation of the finite form of the heterotic Lagrangian under this limit. The main reason of an explicit computation of this Lagrangian is that exact results under the NR limit usually contain large contributions when they are not written in covariant form. We expect that the NR limit of the heterotic supergravity admits a rewriting in terms of $O(D,D+N)$ multiplets. T-duality is an exact symmetry of string theory and the low energy limit of the theory should be able to written in terms of a generalized metric and dilaton. In other words, we expect that the formalism of Double Field Theory [23]-[26] could be able to rewrite the theory, when the fundamental degrees of freedom are parametrized using a non-Riemannian ansatz [27]-[30].
On the other hand, we expect that the vielbein remains gauge invariant while the $\hat{B}$-field inherits a non-covariant gauge transformation (Green-Schwarz mechanism), similarly to what happens in the relativistic case, i.e.,
$\delta_{\hat{\lambda}}\hat{A}_{\mu}^{i}$ $=$ $\partial_{\mu}\hat{\lambda}^{i}+\hat{f}^{i}{}_{jk}\hat{\lambda}^{j}\hat{A}_{\mu}^{k}\,,$ (1.9)
$\delta_{\hat{\lambda}}\hat{B}_{\mu\nu}$ $=$ $-(\partial_{[\mu}\hat{\lambda}^{i})\hat{A}_{\nu]i}\,,$ (1.10)
$\delta_{\hat{\lambda}}\hat{g}_{\mu\nu}$ $=$ $0\,,$ (1.11)
where $\hat{\lambda}$ is a gauge parameter. Due to the previous restrictions, we will expand the gauge field following the proposal given in [22]
$\hat{A}_{\mu}^{i}$ $=c\ {\tau_{\mu}}^{-}\alpha_{-}^{i}+\frac{1}{c}a_{\mu}^{i}\,,$ (1.12)
where $v^{\pm}=\frac{1}{\sqrt{2}}(v^{0}\pm v^{1})$. The expansion (1.12) was constructed in such a way that the fundamental degrees of freedom of DFT remains convergent after taking the NR limit in the double geometry. In this way, we are sure that the NR heterotic supergravity can be rewritten in terms of $O(D,D+N)$ multiplets.
The main goal of this work is to explicitly prove that, when considering $c\rightarrow\infty$, the bosonic part of the heterotic Lagrangian,
$S_{\rm{het}}=\int d^{10}x\sqrt{-\hat{g}}e^{-2\hat{\phi}}(\hat{R}+4\partial_{\mu}\hat{\phi}\partial^{\mu}\hat{\phi}-\frac{1}{12}\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho}-\frac{1}{4}\hat{F}_{\mu\nu}{}^{i}\hat{F}^{\mu\nu}{}_{i})\,,$ (1.13)
with
$\bar{H}_{\mu\nu\rho}=3\left(\partial_{[\mu}\hat{B}_{\nu\rho]}-\hat{C}_{\mu\nu\rho}^{(g)}\right)\,,$ (1.14)
$\hat{C}_{\mu\nu\rho}^{(g)}=\hat{A}^{i}_{[\mu}\partial_{\nu}\hat{A}_{\rho]i}-\frac{1}{3}\hat{f}_{ijk}\hat{A}_{\mu}^{i}\hat{A}_{\nu}^{j}\hat{A}_{\rho}^{k}\,,$ (1.15)
$\hat{F}_{\mu\nu}{}^{i}=2\partial_{[\mu}\hat{A}_{\nu]}{}^{i}-\hat{f}^{i}{}_{jk}\hat{A}_{\mu}{}^{j}\hat{A}_{\nu}{}^{k}\,,$ (1.16)
remains finite. We will also present the explicit form of the gauge transformations and the heteorotic supergravity Lagrangian in its covariant form.
1.1 Organization of the paper
This paper is organized as follows: In the next section we review the cancellation of divergences in bosonic supergravity and we show the cancellation of divergencies in heterotic supergravity. In section 3 we study the gauge transformation rules of the fundamental fields of the theory upon considering the NR limit and we construct a suitable covariant derivative and curvatures for all the fields that transform in a non-covariant fashion. In section 4 we compute the finite Lagrangian of the theory. We will write not only the gauge part using new curvature for the gauge fields, but also the ungauged part of the Lagrangian [13] using new covariant structures which explicit show the invariance under infinitesimal diffeomorphisms. Finally, in section 5, we present a discussion and an outlook.
2 Cancellation of divergences
When one expands the action (1.7), both the $\hat{R}$ and $-\frac{1}{12}\hat{H}^{2}$ terms produce $c^{2}$ contributions. The same happens when we analyze (1.13), but this time the gauge compensation of $-\frac{1}{12}\bar{H}^{2}$ is given also by the $-\frac{1}{4}\hat{F}^{2}$ term. We will analyze both scenarios separately.
2.1 Ungauged case (review)
Let’s review the cancellation of divergences in the bosonic supergravity Lagrangian. To do so, we define
$\hat{R}^{\rho}{}_{\mu\nu\sigma}=-2\partial_{[\mu}\Gamma^{\rho}_{\nu]\sigma}-2\Gamma^{\rho}_{[\mu|\lambda}\Gamma^{\lambda}_{|\nu]\sigma}$ (2.1)
and $\hat{R}=\hat{R}^{\rho}{}_{\mu\nu\sigma}\delta_{\rho}^{\nu}\hat{g}^{\mu\sigma}$. The divergent part of the Ricci scalar, after imposing the Newton-Cartan relations (1.6), is given by
$\hat{R}^{(2)}=-\frac{c^{2}}{2}\eta_{ab}\partial_{\mu}{\tau_{\nu}{}^{a}}\partial_{\rho}{\tau_{\sigma}{}^{b}}(h^{\mu\sigma}h^{\nu\rho}-h^{\mu\rho}h^{\nu\sigma})\,.$ (2.2)
So far we have just considered the expansion of the metric tensor (1.1) and its inverse (1.2). Therefore, for backgrounds where $h^{\mu\nu}$ is constant, the divergence in the Ricci scalar becomes trivial and it is possible to define a non-relativistic supergravity without including a non-relativistic expansion in the $\hat{B}$-field. For a more general scenario we impose (1.3) and then the divergent part of $-\frac{1}{12}\hat{H}^{2}$ is given by
$-\frac{1}{12}(\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho})^{(2)}=-\frac{c^{2}}{2}\epsilon_{ab}\epsilon_{cd}\eta^{ac}\partial_{\mu}\tau_{\nu}{}^{b}\partial_{\rho}\tau_{\sigma}{}^{d}(h^{\mu\rho}h^{\nu\sigma}-h^{\mu\sigma}h^{\nu\rho})\,.$ (2.3)
Using the relation $\epsilon_{ab}\epsilon_{cd}\eta^{ac}=\eta_{bd}$ one can cancel the divergences coming from the Ricci scalar.
2.2 Gauged case
The incorporation of the gauge field $\hat{A}_{\mu}^{i}$ brings new divergences to the supergravity Lagrangian. On the one hand, the Yang-Mills Lagrangian,
$L_{\textrm{YM}}=-\frac{1}{4}\hat{F}_{\mu\nu}^{i}\hat{F}^{\mu\nu}_{i}$ (2.4)
upon considering the non-relativistic expansions (1.12) and (1.2) contains a divergence of the form (see appendix 5 for intermediate steps)
$L_{\textrm{YM}}^{(2)}=-\frac{c^{2}}{2}\alpha_{-}^{i}\alpha_{-i}\partial_{\mu}{\tau_{\nu}^{-}}\partial_{\rho}{\tau_{\sigma}^{-}}(h^{\mu\rho}h^{\nu\sigma}-h^{\mu\sigma}h^{\nu\rho})\,.$ (2.5)
These terms indicate that the $A$-expansion in (1.12) does not lead to a finite Yang-Mills theory. However, the inclusion of Chern-Simons terms in the heterotic Lagrangian (1.13) compensates the divergence (2.5). Indeed, one can compute (we use CADABRA [31] for performing this test. The details are given in the appendix),
$-\frac{1}{12}(\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho})^{(2)}=-R^{(2)}+\frac{1}{4}(\hat{F}_{\mu\nu}^{i}\hat{F}^{\mu\nu}_{i})^{(2)}\,.$ (2.6)
At this point, we have proved that the Heterotic supergravity Lagrangian is finite, but trying to write it in non-covariant form leads to huge number of terms. For this reason, our strategy will be to write it in covariant form, making use of the diffeomorphism and gauge invariance of the action.
In the next section, we will study the form of the gauge transformation rules and we will construct a suitable covariant derivative. After that we will construct new curvatures for all the fundamental fields transforming in a non-covariant way under gauge transformations.
3 Transformations and curvatures
In this section we will focus on two symmetries: infinitesimal diffeomorphism and the non-Abelian gauge transformations associated to the $A$-field.
3.1 Covariant derivative
The covariant derivative of $\tau_{\nu}{}^{a}$ can be easily computed,
$\nabla_{\mu}\tau_{\nu}{}^{a}$ $=$ $\partial_{\mu}\tau_{\nu}{}^{a}-\Gamma_{\mu\nu}^{\rho}\tau_{\rho}{}^{a}$
$=$ $\partial_{\mu}\tau_{\nu}{}^{a}+\partial_{(\mu}{\tau_{\nu)}{}^{a}}-\frac{1}{c^{2}}\partial_{(\mu}{\tau_{\nu)}{}^{a}}+T_{\mu\nu}^{a}\,,$
where $T_{\mu\nu}{}^{a}$ is given by the following covariant quantity,
$T_{\mu\nu}{}^{a}$ $=$ $\eta_{bc}\eta^{ad}\nabla_{(\mu|}{\tau_{\rho}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}-\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu|}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}$ (3.2)
$-\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{(\mu}{\tau_{\rho}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}+\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}\,.$
An interesting observation of $\nabla_{\mu}\tau_{\nu}{}^{a}$ is that this quantity does not diverge, but contains $\frac{1}{c^{2}}$ that we need to track in order to correctly write the finite Lagrangian. This means that the finite Lagrangian will evolve terms with powers of $c$, but this is just the price to pay for writing it in covariant form. The antisymmetric part of $\partial_{\mu}\tau_{\nu}^{a}$ satisfies
$\nabla_{[\mu}\tau_{\nu]}^{a}$ $=$ $\partial_{[\mu}\tau_{\nu]}{}^{a}$ (3.3)
while the symmetric part can be written as
$\partial_{(\mu}\tau_{\nu)}{}^{a}=\frac{c^{2}}{2c^{2}-1}\Big{(}\nabla_{(\mu}\tau_{\nu)}{}^{a}-T_{\mu\nu}{}^{a}\Big{)}=Q_{\mu\nu}{}^{a}\,.$ (3.4)
On the other hand the covariant derivative of $\tau^{\nu}{}_{a}$ is given by
$\nabla_{\mu}\tau^{\nu}{}_{a}=\partial_{\mu}\tau^{\nu}{}_{a}+\Gamma_{\mu\rho}^{\nu}\tau^{\rho}{}_{a}=\frac{1}{2}\partial_{\mu}\tau^{\nu}{}_{a}+S_{\mu}{}^{\nu}{}_{a}\,,$ (3.5)
where $S_{\mu}{}^{\nu}{}_{a}$ is a covariant quantity given by
$S_{\mu}{}^{\nu}{}_{a}$ $=$ $c^{2}\Big{(}\nabla_{[\rho}{\tau_{\sigma]}^{b}}\eta_{bc}\tau_{\mu}^{c}\tau^{\rho}_{a}h^{\nu\sigma}+\nabla_{[\mu}{\tau_{\rho]}^{b}}\eta_{ab}h^{\nu\rho}\Big{)}+\frac{1}{2}Q_{\mu\rho}{}^{b}(\tau^{\nu}_{b}\tau^{\rho}_{a}+\eta_{ab}h^{\nu\rho})$ (3.6)
$+\frac{1}{2}\nabla_{[\mu}{\tau_{\rho]}^{b}}(2\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d}-\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}h^{\nu\rho})+\nabla_{[\rho}{\tau_{\sigma]}^{b}}(\eta_{bc}\tau_{\mu}^{c}\tau^{\sigma}_{a}h^{\nu\rho}+\eta_{bc}\eta^{de}\tau_{\mu}^{c}\tau^{\nu}_{d}\tau^{\rho}_{a}\tau^{\sigma}_{e})$
$+\frac{1}{4c^{2}}\Big{(}Q_{\mu\rho}{}^{b}(-2\tau^{\nu}_{b}\tau^{\rho}_{a}+2\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})+\nabla_{\mu}{\tau_{\rho}^{b}}(\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})$
$-\nabla_{\rho}{\tau_{\mu}^{b}}(\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})+4\nabla_{[\rho}{\tau_{\sigma]}^{b}}\eta_{bc}\eta^{de}\tau_{\mu}^{c}\tau^{\nu}_{d}\tau^{\sigma}_{a}\tau^{\rho}_{e}\Big{)}\,.$
Considering the Newton-Cartan relation, it is easy to prove that the ungauged Lagrangian can be written in terms of $\nabla_{\mu}\tau_{\nu}{}^{a}$, $\nabla_{\mu}\tau^{\nu}{}_{a}$ and the tensors $Q_{\mu\nu}{}^{a}$ and $S_{\mu}{}^{\nu}{}_{a}$, which also depends on the previous covariant derivatives. In other words, we do not need to compute the covariant derivative of $h_{\mu\nu}$ and $h^{\mu\nu}$ to write, for example, the Riemann tensor, the Ricci tensor or the Ricci scalar in covariant form.
3.2 Gauge transformations
Transformation rules under the non-Abelian gauge symmetry can be extracted from (1.9), (1.10) and (1.11). We will impose an extra expansion on the gauge parameter and the structure constants following [22],
$\hat{\lambda}^{i}$ $=$ $\frac{1}{c}\lambda^{i}$ (3.7)
$\hat{f}_{ijk}$ $=$ $cf_{ijk}\,.$ (3.8)
At this point, one can think that the rescaling of the structure constants might include divergences in the gauge Lagrangian. However, this reescaling is safe because higher order $\hat{f}_{ijk}$ contributions are $\propto\frac{1}{c}$. Then, the rescaling only introduces cubic contributions in gauge fields in the finite Lagrangian, which we will present in the next section.
After exploring the gauge transformation (1.9), (1.10) and (1.11) we find the following rules upon taking the limit $c\rightarrow\infty$,
$\delta_{\lambda}a_{\mu}^{i}$ $=$ $\partial_{\mu}\lambda^{i}+f^{i}{}_{jk}\lambda^{j}a_{\mu}^{k}\,,$ (3.9)
$\delta_{\lambda}\alpha_{-}^{i}$ $=$ $f^{i}{}_{jk}\lambda^{j}\alpha_{-}^{k}\,,$ (3.10)
$\delta_{\lambda}b_{\mu\nu}$ $=$ $-(\partial_{[\mu}\lambda^{i})\alpha_{-i}\tau_{\nu]}^{-}\,,$ (3.11)
$\delta_{\lambda}h_{\mu\nu}$ $=$ $0\,,$ (3.12)
$\delta_{\lambda}h^{\mu\nu}$ $=$ $0\,,$ (3.13)
$\delta_{\lambda}\tau_{\mu}{}^{a}$ $=$ $0\,,$ (3.14)
$\delta_{\lambda}\tau^{\mu}{}_{a}$ $=$ $0\,.$ (3.15)
From the previous transformations, we learn that $a_{\mu}^{i}$ transforms as a gauge connection, $\alpha_{-}^{i}$ transforms as a vector and $b_{\mu\nu}$ transforms in a non-covariant way, emulating a Green-Schwarz mechanism. Interestingly, this mechanism cannot be eliminated through field redefinitions in the relativistic scenario. However, when one considers the NR limit, it is possible to define the field
$\bar{b}_{\mu\nu}=b_{\mu\nu}+a_{[\mu|i}\alpha_{-}^{i}\tau_{|\nu]-}$ (3.16)
which is gauge invariant, i.e. $\delta_{\lambda}\bar{b}_{\mu\nu}=0$. This is an interesting property of the heterotic supergravity under the NR limit which will help us to construct a suitable curvature for the $\bar{b}$-field.
3.3 Gauging the covariant derivative
Considering an arbitrary gauge vector $v^{i}$ upon taking the NR limit, the partial derivative acting on this vector is not covariant. Therefore we need to define a covariant derivative for the gauge symmetry in the following way,
$\nabla_{\mu}v^{i}=\partial_{\mu}v^{i}-f^{i}{}_{jk}a_{\mu}{}^{j}v^{k}\,.$ (3.17)
We choose $a_{\mu}{}^{j}$ over $\alpha_{-}^{i}$ to construct the covariant derivative due to the transformation property of $a_{\mu}{}^{j}$. In this case, we need a proper connection and $\alpha_{-}^{i}$ transforms as a gauge vector.
Let’s start constructing the curvature for the gauge connection $a_{\mu}^{i}$, which is given by
$f_{\mu\nu}{}^{i}=2\partial_{[\mu}a_{\nu]}{}^{i}-f^{i}{}_{jk}a_{\mu}{}^{j}a_{\nu}{}^{k}\,,$ (3.18)
The field $\alpha_{-}^{i}$ is already covariant, so we do not need to construct its curvature. If we consider $\bar{b}$ as the fundamental field of the theory, then we define its curvature as
$\bar{h}_{\mu\nu\rho}=3\partial_{[\mu}\bar{b}_{\nu\rho]}\,=h_{\mu\nu\rho}-3c^{(g)}_{\mu\nu\rho}.$ (3.19)
From the previous curvature, we can identify the equivalent quantity to the Chern-Simons terms in this limit,
$c^{(g)}_{\mu\nu\rho}=-\partial_{[\mu}(a_{\nu}^{i}\tau_{\rho]}^{-}\alpha_{-i}).$ (3.20)
In the next section, we will use these curvatures to present the finite NR gauge Lagrangian.
4 Finite Lagrangian
In this section we construct all the finite contributions to the heterotic supergravity Lagrangian in covariant form upon taking the NR limit. For the readers interested only in the final result, we report the final form of the gauge Lagrangian in the conclusions.
4.1 Ungauged Lagrangian
We would like to present the finite heterotic Lagrangian in terms of covariant quantities. We will choose objects which are covariant under infinitesimal diffeomorphisms. Due to the previous section, we have all the pieces to write the ungauged part of the finite Lagrangian in its covariant form. This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term and the H-term given by
$L$ $=$ $R^{(0)}+4\partial_{\mu}\varphi\partial^{\mu}\varphi-\frac{1}{12}H^{(0)}_{\mu\nu\rho}H^{(0)}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$ (4.1)
$-\frac{1}{2}H^{(0)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}H^{(2)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\eta^{cd}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c}\tau^{\sigma}{}_{b}\tau^{\gamma}{}_{d}h^{\rho\epsilon}\,,$
where
$H^{(0)}_{\mu\nu\rho}$ $=$ $h_{\mu\nu\rho}=3\partial_{[\mu}b_{\nu\rho]}\,,$ (4.2)
$H^{(2)}_{\mu\nu\rho}$ $=$ $-6c^{2}\epsilon_{ab}\nabla_{[\mu}\tau_{\nu}{}^{a}\tau_{\rho]}{}^{b}\,,$ (4.3)
and the Ricci scalar is given by
$R^{(0)}$ $=$ $8\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau^{\mu}{}_{a}}h^{\nu\rho}+5\nabla_{[\mu}{\tau_{\nu]}{}^{c}}\eta^{ab}(\nabla_{\rho}{\tau^{\mu}{}_{a}}\tau^{\nu}{}_{c}\tau^{\rho}{}_{b}+\nabla_{\rho}{\tau^{\rho}{}_{a}}\tau^{\mu}{}_{b}\tau^{\nu}{}_{c})$ (4.4)
$-\frac{3}{8}\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}(\tau^{\nu}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\rho}-2\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\rho}+\tau^{\mu}{}_{a}\tau^{\rho}{}_{b}h^{\nu\sigma})$
$+\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}h^{\mu\rho}h^{\nu\sigma}+4c^{2}\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}h^{\mu\sigma}h^{\nu\rho}$
$+\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}\eta^{cd}(2\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}-\tau^{\mu}{}_{c}\tau^{\rho}{}_{d}h^{\nu\sigma}-\tau^{\nu}{}_{c}\tau^{\sigma}{}_{d}h^{\mu\rho})$
$+\frac{1}{c^{2}}\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}\eta^{cd}(\tau^{\nu}{}_{c}\tau^{\sigma}{}_{d}h^{\mu\rho}-2\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}+\tau^{\mu}{}_{c}\tau^{\rho}{}_{d}h^{\nu\sigma})$
$-2\nabla_{\mu}{\nabla_{\nu}{\tau_{\rho}{}^{a}}}\tau^{[\rho}{}_{a}h^{\nu]\mu}+f(Q,S)\,,$
where $f(Q,S)$ is a covariant contribution since both $Q_{\mu\nu}{}^{a}$ and $S_{\mu}{}^{\nu}{}_{a}$ are covariant,
$f(Q,S)$ $=$ $8Q_{\mu\nu}{}^{a}(S_{\rho}{}^{[\mu}{}_{a}h^{\rho]\nu}+\nabla_{\rho}{\tau^{[\rho}{}_{a}}h^{\mu]\nu})-8\nabla_{[\nu}{\tau_{\rho]}{}^{a}}S_{\mu}{}^{\nu}{}_{a}h^{\mu\rho}-\frac{1}{2}Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\tau^{\mu}{}_{a}\tau^{\rho}{}_{b}h^{\nu\sigma}$ (4.5)
$+Q_{\mu\nu}{}^{a}\eta^{bc}(-2\nabla_{\rho}{\tau^{\mu}{}_{b}}\tau^{\nu}{}_{c}\tau^{\rho}{}_{a}+\nabla_{\rho}{\tau^{\mu}{}_{b}}\tau^{\nu}{}_{a}\tau^{\rho}{}_{c}+\nabla_{\rho}{\tau^{\rho}{}_{b}}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c})$
$+Q_{\mu\nu}{}^{a}\eta^{bc}(2S_{\rho}{}^{\mu}{}_{b}\tau^{\nu}{}_{c}\tau^{\rho}{}_{a}-S_{\rho}{}^{\mu}{}_{b}\tau^{\nu}{}_{a}\tau^{\rho}{}_{c}-S_{\rho}{}^{\rho}{}_{b}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c})$
$+\nabla_{\rho}{\tau_{\sigma}{}^{b}}Q_{\mu\nu}{}^{a}(-\frac{3}{2}\tau^{\rho}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\nu}+2\tau^{\mu}{}_{a}\tau^{[\rho}{}_{b}h^{\sigma]\nu}+2\tau^{\rho}{}_{b}\tau^{\sigma}{}_{a}h^{\mu\nu}+4\tau^{\mu}{}_{b}\tau^{[\rho}{}_{a}h^{\sigma]\nu})$
$+\nabla_{\nu}{\tau_{\rho}{}^{c}}\eta^{ab}(-\frac{5}{2}S_{\mu}{}^{\nu}{}_{a}\tau^{\mu}{}_{b}\tau^{\rho}{}_{c}+5S_{\mu}{}^{\mu}{}_{a}\tau^{[\nu}{}_{c}\tau^{\rho]}{}_{b})$
$+\frac{5}{2}\nabla_{\rho}{\tau_{\nu}{}^{c}}\eta^{ab}S_{\mu}{}^{\nu}{}_{a}\tau^{\mu}{}_{b}\tau^{\rho}{}_{c}+2Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\eta_{ab}h^{\mu[\nu}h^{\rho]\sigma}+2Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\eta_{ab}\eta^{cd}\tau^{\mu}{}_{c}\tau^{[\rho}{}_{d}h^{\nu]\sigma}$
$+3\nabla_{\rho}{\tau_{\sigma}{}^{b}}Q_{\mu\nu}{}^{a}\eta_{ab}\eta^{cd}\tau^{\mu}{}_{c}\tau^{[\sigma}{}_{d}h^{\rho]\nu})+\frac{4}{c^{2}}\nabla_{[\rho}{\tau_{\sigma]}{}^{b}}Q_{\mu\nu}{}^{a}(\tau^{\rho}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\nu}+\tau^{\mu}{}_{b}\tau^{\sigma}{}_{a}h^{\nu\rho})$
$+4\nabla_{\mu}{Q_{\nu\rho}{}^{a}}\tau^{[\mu}{}_{a}h^{\nu]\rho}\,.$
In $R^{(0)}$ we can explicitly see that terms proportional to $c^{2}$ and $\frac{1}{c^{2}}$ naturally arises when we write this object using covariant derivatives. When we replace the covariant derivatives by their definitions, all these divergent and convergent quantities cancel, giving a finite contribution in terms of partial derivatives.
4.2 Gauged terms
Before analyzing the gauge contributions of the heteorotic Lagrangian, let’s focus on the ungauged terms depending on the b-field,
$L_{b}$ $=$ $-\frac{1}{4}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{b_{\gamma\epsilon}}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\sigma}{}_{c}h^{\mu\rho}h^{\nu\gamma}+\frac{1}{2}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{b_{\gamma\epsilon}}h^{\mu\gamma}h^{\nu\sigma}h^{\rho\epsilon}$ (4.6)
$-\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\rho}{}_{c}h^{\mu\sigma}h^{\nu\gamma}-\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}\,.$
These terms are not gauge invariant, due to the transformation (3.11). Therefore, we rewrite these contributions in terms of the new curvature for the b-field,
$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}{}^{b}\bar{h}_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}+\nabla_{\mu}\tau_{\nu}^{-}f_{\gamma\rho}{}^{i}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}+L_{\textrm{comp}}\,,$ (4.7)
where $L_{\textrm{comp}}=L^{(2)}_{\textrm{non-cov}}+L^{(3)}_{\textrm{non-cov}}+L^{(4)}_{\textrm{non-cov}}$ with
$L^{(2)}_{\textrm{non-cov}}$ $=$ $\nabla_{\mu}\tau_{\nu}^{-}a_{[\rho}^{i}\nabla_{\gamma]}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}+3\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}a_{[\sigma}^{i}\alpha_{-i}\nabla_{\rho}\tau_{\gamma]}^{-}\tau^{\sigma}_{c}h^{\mu\gamma}h^{\nu\rho}$ (4.8)
$+\frac{1}{2}h_{\mu\nu\rho}a_{\gamma}^{i}\alpha_{-i}\nabla_{\sigma}(\tau_{\epsilon}^{-})h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}\,,$
$L^{(3)}_{\textrm{non-cov}}=-\nabla_{\mu}\tau_{\nu}^{-}f^{i}{}_{jk}a_{\gamma}^{j}a_{\rho}{}^{k}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}\,,$ (4.9)
and
$L^{(4)}_{\textrm{non-cov}}$ $=$ $\frac{3}{4}a_{[\nu}^{i}\alpha_{-i}\nabla_{\mu}(\tau_{\rho]^{-}})a_{\gamma}^{j}\alpha_{-j}\nabla_{\sigma}\tau_{\epsilon}^{-}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}\,.$ (4.10)
In the next subsection we will study the contributions to the gauge Lagrangian coming from the Yang-Mills terms and the Chern-Simons contributions. We split the presentation according to an expansion in gauge fields.
4.3 Quadratic gauged contributions
4.3.1 Quadratic Yang-Mills contributions
The quadratic contributions coming from $-\frac{1}{4}\hat{F}^{2}$ are given by
$-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$
$+2\alpha_{-i}\eta^{ab}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau_{\sigma}^{-}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-2\alpha_{-i}\nabla_{\mu}\tau_{\nu}{}^{-}\nabla_{[\rho}a_{\sigma]}^{i}h^{\mu\rho}h^{\nu\sigma}\,.$ (4.11)
The previous expression can be written in a covariant form as
$-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$
$-2\alpha_{-i}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau^{\nu}_{+}h^{\mu\rho}-\alpha_{-i}\nabla_{\mu}\tau_{\nu}{}^{-}f_{\rho\sigma}^{i}h^{\mu\rho}h^{\nu\sigma}+L^{(3)}(a,\alpha)$ (4.12)
where the inclusion of the (gauge) covariant derivative and the curvature $f_{\mu\nu i}$ introduces cubic contributions
$L^{(3)}(a,\alpha)=-\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}f^{i}{}_{jk}a^{j}_{\rho}a^{k}_{\sigma}h^{\mu\rho}h^{\nu\sigma}+2\alpha_{-i}\eta^{ab}f^{i}{}_{jk}a_{\mu}^{j}\alpha_{-}^{k}\nabla_{[\nu}\tau_{\rho]}^{-}\tau_{\sigma}^{-}\tau_{a}^{\nu}\tau_{b}^{\sigma}h^{\mu\rho}\,.$ (4.13)
4.3.2 Quadratic Chern-Simons contributions
The quadratic contributions coming from $-\frac{1}{12}\bar{H}^{2}$ are given by
$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}$
$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}$
$-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}a_{\sigma]i}h^{\mu\sigma}h^{\nu\rho}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}h_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}-L^{(2)}_{\textrm{non-cov}}\,.$ (4.14)
Among these terms we have identified all the quadratic non-covariant contributions needed in (4.7). We can write the previous expression in covariant form as
$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}$
$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}$
$-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}f_{\rho\sigma i}h^{\mu\sigma}h^{\nu\rho}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}\bar{h}_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}-L^{(2)}_{\textrm{non-cov}}+\tilde{L}^{(3,4)}(a,\alpha)\,,$ (4.15)
where once again the inclusion of the curvatures $f_{\mu\nu i}$ and $\bar{h}_{\rho\sigma\gamma}$ in the last line of the previous expression introduces cubic and quartic contributions, respectively, that we will take into account when we analyze the next orders. We capture all these new contributions under the notation $\tilde{L}^{(3,4)}(a,\alpha)$, which splits in the following form
$\tilde{L}^{(3)}(a,\alpha)=-\alpha_{-}^{i}\nabla_{\mu}\tau_{\nu}^{-}f_{i}{}^{jk}a_{\rho j}a_{\sigma k}h^{\mu\sigma}h^{\nu\rho}\,,$ (4.16)
and
$\tilde{L}^{(4)}(a,\alpha)=\frac{3}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}a_{[\sigma}^{j}\alpha_{-j}\nabla_{\rho}(\tau_{\gamma]}^{-})\tau_{+}^{\rho}h^{\mu\sigma}h^{\nu\gamma}\,.$ (4.17)
4.4 Cubic gauge contributions
4.4.1 Cubic Yang-Mills contributions
The cubic contributions coming from $-\frac{1}{4}\hat{F}^{2}$ are given by
$-L^{(3)}_{\textrm{non-cov}}-L^{(3)}(a,\alpha)-\tilde{L}^{(3)}(a,\alpha)\,.$ (4.18)
The previous expression exactly cancels the contributions obtained in (4.13) and (4.16) and there are no additional contributions at this order in fields.
4.5 Quartic gauge contributions
4.5.1 Quartic Chern-Simons contributions
The quartic terms in $-\frac{1}{12}\bar{H}^{2}$ are given by
$-\tilde{L}^{(4)}(a,\alpha)\,,$ (4.19)
and there are no other additional contributions at this order, or higher, in fields.
5 Discussion
In this work we have studied the finite form of the heterotic supergravity Lagrangian, extending the formulation of [13]. The Ricci scalar part of the Lagrangian has been rewritten in terms of $\nabla_{\mu}\tau_{\nu}{}^{a}$, $\nabla_{\mu}\tau^{\nu}{}_{a}$, $Q_{\mu\nu}{}^{a}$ (eq. (3.4)) and $S_{\mu}{}^{\nu}{}_{a}$ (eq. (3.6)), showing the explicit diffeomorphism invariance of the theory.
When we include the heteoric gauge field, the bosonic part of the Lagrangian of this theory contains two divergent parts given by the Chern-Simons terms and the Yang-Mills Lagrangian which exactly compensate each other, as it happens in bosonic supergravity, giving rise to a finite Lagrangian. The gauge Lagrangian depends on a pair of gauge fields, $a_{\mu}^{i}$ and $\alpha_{-}^{i}$, the former transforming as a gauge connection, so we constructed its curvature, and the latter transforming as a gauge vector, so we do not need to construct a curvature. Interestingly enough, the Green-Schwarz mechanism for the b-field can be trivialized in this limit and therefore one can define $\tilde{b}_{\mu\nu}$ as in (3.16), and its curvature $\tilde{h}_{\mu\nu\rho}$, both of them gauge invariant. The gauge contributions in $\tilde{h}_{\mu\nu\rho}$ can be though of as Chern-Simons contributions.
In this work we have presented the finite Lagrangian of the theory which contains quadratic, cubic and quartic contributions. The full gauge Lagrangian is given by
$L^{(g)}_{\textrm{het}}$ $=$ $-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$ (5.1)
$-2\alpha_{-i}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau^{\nu}_{+}h^{\mu\rho}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}$
$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}$
$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}f_{\rho\sigma i}h^{\mu\sigma}h^{\nu\rho}$
$-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}\bar{h}_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}\,,$
while the ungauged action must include the Chern-Simons terms in the curvature of the b-field as we show in (4.7), i.e.,
$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}{}^{b}\bar{h}_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}\,.$ (5.2)
5.1 Outlook
As promising avenue for future work, it would be very interesting to explore the incorporation of $\alpha´$-corrections to both the bosonic and the heterotic supergravity  [^2] See [32] for a pedagogical review on this topic and references therein. in the NR limit. Since the Riemann tensor contains divergences, one should study whether these are controlled by the $\hat{H}_{\mu\nu\rho}$ (or $\bar{H}_{\mu\nu\rho}$ in the heterotic case) contributions, as happens at the two derivative level. The covariant way of writing the Ricci scalar using the covariant derivative of $\tau_{\mu}{}^{a}$ and $\tau^{\mu}{}_{a}$ could be easily extended in order to incorporate the spin connection $\omega_{\mu ab}$, so that the action is also invariant under the $SO(1,1)$ transformations. In this way, the higher-derivative contributions could be written in a more compact form, since we expect large contributions if the curvatures are not chosen in an optimal way.
While DFT has been proved a powerful tool to read the expansion of the A-field [22], taking the NR limit at the DFT level including higher-derivative terms is probably the best strategy to read these corrections and then parametrize them to the supergravity level.
Acknowledgements
E.L. is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.
Appendix A Details on the cancellation of divergences: gauge case
The field strength $F_{\mu\nu}{}^{i}$, after expanding $A_{\mu}{}^{i}$ considering (1.12) is given by
$F^{(1)}_{\mu\nu}{}^{i}=\alpha_{-}^{i}\partial_{\mu}{\tau_{\nu}^{-}}+\partial_{\mu}{\alpha_{-}^{i}}\tau_{\nu}^{-}-\alpha_{-}^{i}\partial_{\nu}{\tau_{\mu}^{-}}-\partial_{\nu}{\alpha_{-}^{i}}\tau_{\mu}^{-}\,,$ (A.1)
$F^{(0)}_{\mu\nu}{}^{i}=+\alpha_{-}^{j}\eta^{ik}\tau_{\mu}^{-}a_{\nu}^{i1}f_{jki1}-\alpha_{-}^{j}\eta^{ik}\tau_{\nu}^{-}a_{\mu}^{i1}f_{jki1}\,,$ (A.2)
$F^{(-1)}_{\mu\nu}{}^{i}=\partial_{\mu}{a_{\nu}^{i}}-\partial_{\nu}{a_{\mu}^{i}}\,,$ (A.3)
$F^{(-2)}_{\mu\nu}{}^{i}=-\eta^{ij}a_{\mu}^{k}a_{\nu}^{i1}f_{jki1}\,,$ (A.4)
where the notation $F^{(i)}_{\mu\nu}{}^{i}$ denotes the terms of order $c^{i}$ in the expansion. The contraction of this field strength with two inverse metrics, $F^{\mu\nu i}=F_{\rho\sigma}g^{\mu\rho}g^{\nu\sigma}$ yields the expression,
$F^{\mu\nu i}=F^{(1)\mu\nu i}+\mathcal{O}\big{(}\frac{1}{c}\big{)}\,,$ (A.5)
with
$F^{(1)\mu\nu i}=\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\rho}h^{\nu\sigma}-\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\sigma}h^{\nu\rho}\,.$ (A.6)
As the reader can easily check, the only contraction which produces non-vanishing divergences is given by $F_{(1)\mu\nu}{}^{i}F^{(1)\mu\nu j}\eta_{ij}$, which leads to the result (2.5).
For the $-\frac{1}{12}\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho}$ contributions, an important observation is that
$\bar{H}_{\mu\nu\rho}=\bar{H}^{(2)}_{\mu\nu\rho}+\bar{H}_{(0)\mu\nu\rho}+\mathcal{O}\big{(}\frac{1}{c}\big{)}\,,$ (A.7)
while the same object when all the indices raised by the inverse metric has the following form,
$\bar{H}^{\mu\nu\rho}=\bar{H}^{(0)\mu\nu\rho}+\mathcal{O}\big{(}\frac{1}{c^{2}}\big{)}\,.$ (A.8)
Thus, we focus on the objects that contribute to divergences,
$\bar{H}^{(2)}_{\mu\nu\rho}$ $=$ $-\epsilon_{ab}\partial_{\mu}{\tau_{\nu}^{a}}\tau_{\rho}^{b}+\epsilon_{ab}\partial_{\mu}{\tau_{\rho}^{a}}\tau_{\nu}^{b}-\epsilon_{ab}\partial_{\nu}{\tau_{\rho}^{a}}\tau_{\mu}^{b}+\epsilon_{ab}\partial_{\nu}{\tau_{\mu}^{a}}\tau_{\rho}^{b}-\epsilon_{ab}\partial_{\rho}{\tau_{\mu}^{a}}\tau_{\nu}^{b}$ (A.9)
$+\epsilon_{ab}\partial_{\rho}{\tau_{\nu}^{a}}\tau_{\mu}^{b}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\nu}{\tau_{\rho}^{-}}\tau_{\mu}^{-}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\rho}{\tau_{\mu}^{-}}\tau_{\nu}^{-}$
$-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{-}}\tau_{\rho}^{-}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\rho}{\tau_{\nu}^{-}}\tau_{\mu}^{-}$
$+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\mu}{\tau_{\rho}^{-}}\tau_{\nu}^{-}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}\partial_{\nu}{\tau_{\mu}^{-}}\tau_{\rho}^{-}\,,$
$\bar{H}^{(0)\mu\nu\rho}$ $=$ $h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\rho}_{c}h^{\mu\sigma}h^{\nu\gamma}-\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\nu}_{c}h^{\mu\sigma}h^{\rho\gamma}$ (A.10)
$+\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\mu}_{c}h^{\nu\sigma}h^{\rho\gamma}-\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\rho}_{c}h^{\mu\gamma}h^{\nu\sigma}+\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\nu}_{c}h^{\mu\gamma}h^{\rho\sigma}$
$-\epsilon_{ab}\eta^{ac}\partial_{\sigma}{\tau_{\gamma}^{b}}\tau^{\mu}_{c}h^{\nu\gamma}h^{\rho\sigma}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\mu}_{a}\tau^{\epsilon}_{b}h^{\nu\sigma}h^{\rho\gamma}$
$-\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\epsilon}h^{\nu\sigma}h^{\rho\gamma}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\nu}_{a}\tau^{\epsilon}_{b}h^{\mu\gamma}h^{\rho\sigma}$
$-\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\gamma}h^{\nu\epsilon}h^{\rho\sigma}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\rho}_{a}\tau^{\epsilon}_{b}h^{\mu\sigma}h^{\nu\gamma}$
$-\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\mu}_{a}\tau^{\epsilon}_{b}h^{\nu\gamma}h^{\rho\sigma}$
$+\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\epsilon}h^{\nu\gamma}h^{\rho\sigma}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\nu}_{a}\tau^{\epsilon}_{b}h^{\mu\sigma}h^{\rho\gamma}$
$+\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\sigma}h^{\nu\epsilon}h^{\rho\gamma}+\frac{1}{2}\alpha_{-}^{i}\alpha_{-}^{j}\eta^{ab}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}\tau_{\epsilon}^{-}\tau^{\rho}_{a}\tau^{\epsilon}_{b}h^{\mu\gamma}h^{\nu\sigma}$
$+\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\gamma}h^{\nu\sigma}h^{\rho\epsilon}\,.$
The multiplication of the previous two quantities leads directly to the result (2.6).

### v5

@@
 Abstract
-We show that the $D=10$ heterotic supergravity in a non-relativistic (NR) limit has a finite Lagrangian due to non-trivial cancellations of divergent parts arising from the Chern-Simons terms in the curvature of the $\hat{B}$-field and the Yang-Mills Lagrangian.
-This is similar to what happens in bosonic supergravity between the Ricci scalar, $\hat{R}$, and the $-\frac{1}{12}\hat{H}^{2}$ term, under the same limit.
-In this work we present the explicit form of the gauge transformations and curvatures after considering a NR limit and we compute the full finite Lagrangian in its covariant form.
-Particularly, we address both the ungauged part and the gauged part, and we write it using covariant derivatives and curvatures for the gauge fields.
-These results constitutes the first computation of a finite Lagrangian for heterotic supergravity under these kinds of limits.
-As an interesting feature of the expansion implemented, the Green-Schwarz mechanism for the two-form can be trivialized and terms equivalent to Chern-Simons contributions naturally arise from the previous property.
+We show that the $D=10$ heterotic supergravity under a non-relativistic (NR) limit has a finite Lagrangian due to non-trivial cancellations of divergent parts arising from the Chern-Simons terms in the curvature of the $\hat{B}-$field and the Yang-Mills Lagrangian.
+This is similar to what happens in bosonic supergravity between the Ricci scalar, $\hat{R}$, and the $\frac{1}{12}\hat{H}^{2}$ term, under the same limit.
+In this work we present the explicit form of the gauge transformations and gauge curvatures after considering a NR limit.
+We compute the full finite Lagrangian in its gauge covariant form, using gauge covariant curvatures and gauge covariant derivatives.
+These results constitute the first computation of a finite Lagrangian for heterotic supergravity under these kinds of limits.
+As an interesting feature of the expansion implemented, the Green-Schwarz mechanism for the two-form can be trivialized.
+Terms equivalent to Chern-Simons contributions naturally arise from the previous property.
 1 Introduction
-In the last decade there has been huge progress made in the study of the different formulations of non-relativistic string theory [1]-[3] and its supergravity limit [4]-[17]  [^1] See [18]-[20] for reviews on this topic..
-The low energy limit of bosonic string theory is given by the universal NS-NS supergravity, whose fundamental fields $\hat{g}$, $\hat{B}$, $\hat{\phi}$ admit a non-relativistic expansion of the form [13],
+In the last decade, there has been significant progress made in the study of the different formulations of non-relativistic string theory [1]-[3] and its supergravity limit [4]-[17]  [^1] See [18]-[19] for reviews, and also [20] for a complementary introduction to this topic..
+The low-energy limit of bosonic string theory is given by the universal NS-NS supergravity, whose fundamental fields $\hat{g}$, $\hat{B}$, $\hat{\phi}$ admit a non-relativistic expansion of the form [13],
 $\hat{g}_{\mu\nu}$ $=$ $c^{2}{\tau_{\mu}}^{a}{\tau_{\nu}}^{b}\eta_{ab}+h_{\mu\nu}\,$, (1.1)
@@
 $\hat{\phi}$ $=$ $\ln(c)+\varphi\,$, (1.4)
-where $\mu,\nu=1,\dots,10$ are space-time indices and we have split the flat index $\hat{a}=(a,a^{\prime})$ where $a,b,...=0,1$ are the transverse directions and $a^{\prime},b^{\prime},...=2,\dots,9$.
+where $\mu,\nu=1,\dots,10$ are space-time indices and we have split the flat index $\hat{a}=(a,a^{\prime})$ where a,b,=0,1a,b,...=0,1 are the longitudinal directions and $a^{\prime},b^{\prime},...=2,\dots,9$.
 The quantities in the previous expansion obey the following Newton-Cartan relations,
@@
 $\tau_{\mu}{}^{a}\tau^{\mu b}$ $=\eta^{ab}\,,\qquad\tau_{\mu}{}^{a}\tau^{\nu}{}_{a}+e_{\mu}{}^{a^{\prime}}e^{\nu}{}_{a^{\prime}}=\delta_{\mu}^{\nu}\,.$ (1.6)
-While, apriori, this expansion produces divergent tensorial quantities, i.e., Riemann tensor and its traces, H strength, etc., the NS-NS supergravity Lagrangian,
+While, a priori, this expansion produces divergent tensorial quantities upon taking the NR limit $c\rightarrow\infty$ (i.e.
+Riemann tensor and its traces, H strength, etc.), the NS-NS supergravity Lagrangian,
 $S=\int d^{10}x\sqrt{-\hat{g}}e^{-2\hat{\phi}}\Big(\hat{R}+4\partial_{\mu}\hat{\phi}\partial^{\mu}\hat{\phi}-\frac{1}{12}\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho}\Big)\,$, (1.7)
@@
 $\hat{H}_{\mu\nu\rho}=3\partial_{[\mu}\hat{B}_{\nu\rho]}\,$, (1.8)
-remains finite due to a curious cancellation between the divergent part of the Ricci scalar and the divergent part of the $\hat{H}^{2}$ term [13].
+remains finite due to a curious cancellation between the divergent part of the Ricci scalar and the divergent part of the $\hat{H}$ squared term [13].
 In the case of the bosonic sector of heterotic supergravity, we expect a similar cancellation when including the gauge field $\hat{A}_{\mu}^{i}$, $i=1,\dots,N$ and $N=496$ which is the dimension of the heterotic gauge group.
 Given a particular expansion for the gauge field, the inclusion of Chern-Simons terms in the curvature of the $\hat{B}-$field should produce divergences that must be compensated by the Yang-Mills Lagrangian.
-While there are several candidates for the expansion of the A-field in the heterotic scenario [21]-[22], there is still no explicit computation of the finite form of the heterotic Lagrangian under this limit.
-The main reason of an explicit computation of this Lagrangian is that exact results under the NR limit usually contain large contributions when they are not written in covariant form.
-We expect that the NR limit of the heterotic supergravity admits a rewriting in terms of $O(D,D+N)$ multiplets.
-T-duality is an exact symmetry of string theory and the low energy limit of the theory should be able to written in terms of a generalized metric and dilaton.
-In other words, we expect that the formalism of Double Field Theory [23]-[26] could be able to rewrite the theory, when the fundamental degrees of freedom are parametrized using a non-Riemannian ansatz [27]-[30].
-On the other hand, we expect that the vielbein remains gauge invariant while the $\hat{B}$-field inherits a non-covariant gauge transformation (Green-Schwarz mechanism), similarly to what happens in the relativistic case, i.e.,
+A first expansion for the $A$-field was constructed in [21], where the authors showed the cancellation of divergences due to the new gauge contributions.
+Another possible expansion was given in [22], using a manifest T-duality scenario, which also provides a finite supergravity guaranteed by T-duality invariance.
+Despite these recent efforts in this direction, there is still no explicit computation of the finite form of the heterotic Lagrangian under the NR limit, which is one of the goals of this work.
+[…]
+$\hat{\cal H}_{MN}$ $=$ $\hat{\cal H}^{(0)}_{MN}+\frac{1}{c^{2}}\hat{\cal H}^{(-2)}_{MN}\,$ (1.9)
+$e^{-2\hat{d}}$ $=$ $e^{-2\hat{\Phi}}\sqrt{-\hat{g}}=e^{-2\phi}\sqrt{f(\tau,h)}=e^{-2d}\,$, (1.10)
+where $f(\tau,h)=-\frac{\hat{g}}{c^{4}}$ and the different contributions to the generalized metric are given by
+${\cal H}^{(0)}_{MN}$ $=\left(\begin{matrix}h^{\mu\nu}&\epsilon_{ab}\tau_{\nu}{}^{b}\tau^{\mu a}-b_{\rho\nu}h^{\rho\mu}\\ \epsilon_{ab}\tau_{\mu}{}^{b}\tau^{\nu a}-b_{\rho\mu}h^{\rho\nu}&h_{\mu\nu}+b_{\rho\mu}h^{\rho\sigma}b_{\sigma\nu}-2\epsilon_{cd}\tau_{(\mu|}{}^{d}b_{\sigma|\nu)}\tau^{\sigma c}\,\end{matrix}\right)$ (1.11)
+${\cal H}^{(-2)}_{MN}$ $=\left(\begin{matrix}\tau^{\mu a}\tau^{\nu}{}_{a}&-b_{\rho\nu}\tau^{\rho c}\tau^{\mu}{}_{c}\\ -b_{\rho\mu}\tau^{\rho c}\tau^{\nu}{}_{c}&b_{\rho\mu}\tau^{\rho c}b_{\sigma\nu}\tau^{\sigma}{}_{c}\end{matrix}\right)\,.$ (1.12)
+[…]
+The extension from bosonic DFT to heterotic DFT can be easily done by extending the duality group from $O(D,D)$ to $O(D,D+N)$, where $N$ is the dimension of the heterotic gauge group.
+The components of the generalized metric in this case contain a metric $\hat{g}_{\mu\nu}$, a $\hat{B}-$field and a gauge field $\hat{A}_{\mu}{}^{i}$, which transforms as a connection under the gauge group.
+The $\hat{B}-$field inherits a non-covariant gauge transformation (Green-Schwarz mechanism), i.e.,
 $\delta_{\hat{\lambda}}\hat{A}_{\mu}^{i}$ $=$ $\partial_{\mu}\hat{\lambda}^{i}+\hat{f}^{i}{}_{jk}\hat{\lambda}^{j}\hat{A}_{\mu}^{k}\,$, (1.13)
 $\delta_{\hat{\lambda}}\hat{B}_{\mu\nu}$ $=$ $-(\partial_{[\mu}\hat{\lambda}^{i})\hat{A}_{\nu]i}\,\,$, (1.14)
-$\delta_{\hat{\lambda}}\hat{g}_{\mu\nu}$ $=$ $0\,,$ (1.11)
-where $\hat{\lambda}$ is a gauge parameter. Due to the previous restrictions, we will expand the gauge field following the proposal given in [22]
+with $\hat{\lambda}$ an arbitrary gauge parameter.
+In order to preserve the finiteness of the generalized metric when including the gauge field one possibility is to expand the field $\hat{A}_{\mu}{}^{i}$ following the proposal [22]
 $\hat{A}_{\mu}^{i}$ $=c\ {\tau_{\mu}}^{-}\alpha_{-}^{i}+\frac{1}{c}a_{\mu}^{i}\,$, (1.15)
 where $v^{\pm}=\frac{1}{\sqrt{2}}(v^{0}\pm v^{1})$.
-The expansion (1.12) was constructed in such a way that the fundamental degrees of freedom of DFT remains convergent after taking the NR limit in the double geometry.
-In this way, we are sure that the NR heterotic supergravity can be rewritten in terms of $O(D,D+N)$ multiplets.
+The expansion (1.15) was constructed in such a way that the generalized metric of heterotic DFT, which is an element of $O(D,D+N)$, remains convergent after taking the NR limit in the double geometry.
+Demanding this condition from the very beginning allows one to have a finite action and equations of motion, by construction, upon considering the NR limit in the double geometry, since both the Lagrangian and the equations of motion depend on the generalized metric and dilaton (metric formulation of DFT).
+The construction of non-relativistic heterotic supergravity given in this work (which continues the ideas formulated in [22]) is an alternative formulation to the one presented in [21], and both constructions should be related by field redefinitions in order to have a unique supersymmetric extension (see [15] for the minimal supersymmetric extension).
 The main goal of this work is to explicitly prove that, when considering $c\rightarrow\infty$, the bosonic part of the heterotic Lagrangian,
-$S_{\rm{het}}=\int d^{10}x\sqrt{-\hat{g}}e^{-2\hat{\phi}}(\hat{R}+4\partial_{\mu}\hat{\phi}\partial^{\mu}\hat{\phi}-\frac{1}{12}\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho}-\frac{1}{4}\hat{F}_{\mu\nu}{}^{i}\hat{F}^{\mu\nu}{}_{i})\,,$ (1.13)
+[…]
 with
@@
 $\hat{F}_{\mu\nu}{}^{i}=2\partial_{[\mu}\hat{A}_{\nu]}{}^{i}-\hat{f}^{i}{}_{jk}\hat{A}_{\mu}{}^{j}\hat{A}_{\nu}{}^{k}\,$, (1.19)
-remains finite. We will also present the explicit form of the gauge transformations and the heteorotic supergravity Lagrangian in its covariant form.
+remains finite. We will also present the explicit form of the symmetry transformations and the full heterotic supergravity Lagrangian in its gauge covariant form upon considering the NR limit.
 1.1 Organization of the paper
-This paper is organized as follows: In the next section we review the cancellation of divergences in bosonic supergravity and we show the cancellation of divergencies in heterotic supergravity.
-In section 3 we study the gauge transformation rules of the fundamental fields of the theory upon considering the NR limit and we construct a suitable covariant derivative and curvatures for all the fields that transform in a non-covariant fashion.
-In section 4 we compute the finite Lagrangian of the theory.
-We will write not only the gauge part using new curvature for the gauge fields, but also the ungauged part of the Lagrangian [13] using new covariant structures which explicit show the invariance under infinitesimal diffeomorphisms.
+This paper is organized as follows: in the next section we review the cancellation of divergences in bosonic supergravity and we show the cancellation of divergences in heterotic supergravity.
+In section 3 we study the gauge transformation rules of the fundamental fields of the theory upon considering the NR limit and we construct a suitable gauge covariant derivative and gauge curvatures for the gauge fields.
+In section 4 we compute the finite Lagrangian of the theory in its gauge covariant form.
 Finally, in section 5, we present a discussion and an outlook.
 2 Cancellation of divergences
-When one expands the action (1.7), both the $\hat{R}$ and $-\frac{1}{12}\hat{H}^{2}$ terms produce $c^{2}$ contributions.
-The same happens when we analyze (1.13), but this time the gauge compensation of $-\frac{1}{12}\bar{H}^{2}$ is given also by the $-\frac{1}{4}\hat{F}^{2}$ term.
+When one expands the action (1.7), both the $\hat{R}$ and $\frac{1}{12}\hat{H}^{2}$ terms produce $c^{2}$ contributions.
+The same happens when we analyze (4.7), but this time the gauge compensation of $\frac{1}{12}\bar{H}^{2}$ is given also by the $\frac{1}{4}\hat{F}^{2}$ term.
 We will analyze both scenarios separately.
 2.1 Ungauged case (review)
-Let’s review the cancellation of divergences in the bosonic supergravity Lagrangian. To do so, we define
+Let us review the construction given in [13], with a focus on the cancellation of divergences in the bosonic supergravity Lagrangian. To do so, we define
 $\hat{R}^{\rho}{}_{\mu\nu\sigma}=-2\partial_{[\mu}\Gamma^{\rho}_{\nu]\sigma}-2\Gamma^{\rho}_{[\mu|\lambda}\Gamma^{\lambda}_{|\nu]\sigma}$ (2.1)
 and $\hat{R}=\hat{R}^{\rho}{}_{\mu\nu\sigma}\delta_{\rho}^{\nu}\hat{g}^{\mu\sigma}$. The divergent part of the Ricci scalar, after imposing the Newton-Cartan relations (1.6), is given by
-$\hat{R}^{(2)}=-\frac{c^{2}}{2}\eta_{ab}\partial_{\mu}{\tau_{\nu}{}^{a}}\partial_{\rho}{\tau_{\sigma}{}^{b}}(h^{\mu\sigma}h^{\nu\rho}-h^{\mu\rho}h^{\nu\sigma})\,.$ (2.2)
+[…]
 So far we have just considered the expansion of the metric tensor (1.1) and its inverse (1.2).
 Therefore, for backgrounds where $h^{\mu\nu}$ is constant, the divergence in the Ricci scalar becomes trivial and it is possible to define a non-relativistic supergravity without including a non-relativistic expansion in the $\hat{B}-$field.
-For a more general scenario we impose (1.3) and then the divergent part of $-\frac{1}{12}\hat{H}^{2}$ is given by
-$-\frac{1}{12}(\hat{H}_{\mu\nu\rho}\hat{H}^{\mu\nu\rho})^{(2)}=-\frac{c^{2}}{2}\epsilon_{ab}\epsilon_{cd}\eta^{ac}\partial_{\mu}\tau_{\nu}{}^{b}\partial_{\rho}\tau_{\sigma}{}^{d}(h^{\mu\rho}h^{\nu\sigma}-h^{\mu\sigma}h^{\nu\rho})\,.$ (2.3)
-Using the relation $\epsilon_{ab}\epsilon_{cd}\eta^{ac}=\eta_{bd}$ one can cancel the divergences coming from the Ricci scalar.
+For a more general scenario we impose (1.3) and then the divergent part of $\frac{1}{12}\hat{H}^{2}$ is given by
+[…]
+Using the relation $\epsilon_{ab}\epsilon_{cd}\eta^{ac}=\eta_{bd}$ one can cancel the divergences coming from the Ricci scalar term.
 2.2 Gauged case
@@
 $L_{\textrm{YM}}=-\frac{1}{4}\hat{F}_{\mu\nu}^{i}\hat{F}^{\mu\nu}_{i}$ (2.4)
-upon considering the non-relativistic expansions (1.12) and (1.2) contains a divergence of the form (see appendix 5 for intermediate steps)
+upon considering the non-relativistic expansions (1.15) and (1.2), contains a divergence of the form (see appendix A for intermediate steps)
 $L_{\textrm{YM}}^{(2)}=-\frac{c^{2}}{2}\alpha_{-}^{i}\alpha_{-i}\partial_{\mu}{\tau_{\nu}^{-}}\partial_{\rho}{\tau_{\sigma}^{-}}(h^{\mu\rho}h^{\nu\sigma}-h^{\mu\sigma}h^{\nu\rho})\,.$ (2.5)
@@
 However, the inclusion of Chern-Simons terms in the heterotic Lagrangian (4.7) compensates the divergence (2.5).
-Indeed, one can compute (we use CADABRA [31] for performing this test.
-The details are given in the appendix),
+Indeed, one can compute (we have used CADABRA [31] for performing this test, but the main details of the cancellation are given in the Appendix A),
 $-\frac{1}{12}(\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho})^{(2)}=-R^{(2)}+\frac{1}{4}(\hat{F}_{\mu\nu}^{i}\hat{F}^{\mu\nu}_{i})^{(2)}\,.$ (2.6)
-At this point, we have proved that the Heterotic supergravity Lagrangian is finite, but trying to write it in non-covariant form leads to huge number of terms.
-For this reason, our strategy will be to write it in covariant form, making use of the diffeomorphism and gauge invariance of the action.
-In the next section, we will study the form of the gauge transformation rules and we will construct a suitable covariant derivative.
-After that we will construct new curvatures for all the fundamental fields transforming in a non-covariant way under gauge transformations.
+At this point, we have proved that the heterotic supergravity Lagrangian is finite, but trying to write it in a non-covariant form leads to a huge number of terms.
+For this reason, our strategy will be to write it in a gauge covariant form, making use of the gauge invariance of the action.
+In the next section, we will study the form of the gauge transformation rules and we will construct a suitable gauge covariant derivative.
+After that, we will construct new curvatures for all the fundamental fields transforming in a non-covariant way under gauge transformations.
 3 Transformations and curvatures
-In this section we will focus on two symmetries: infinitesimal diffeomorphism and the non-Abelian gauge transformations associated to the $A$-field.
+In this section we will focus on the non-Abelian gauge transformations, in order to be able to write a gauge covariant action using tensors under this symmetry.
+We will start by describing the Levi-Civita connection, then we will construct the gauge covariant derivative and finally we will include gauge transformations with a gauging in the covariant derivative.
+Remarkably, we will describe the Green-Schwarz mechanism for the b-field and its trivialization, considering field redefinitions.
+3.1 Levi-Civita connection
+The Levi-Civita connection can be written as
+$\Gamma_{\mu\nu}^{\rho}=c^{2}\Gamma_{\mu\nu}^{(2)\rho}+\Gamma_{\mu\nu}^{(0)\rho}+\frac{1}{c^{2}}\Gamma_{\mu\nu}^{(-2)\rho}\,$, (3.1)
+where the explicit form of the previous terms is given by
+$\Gamma_{\mu\nu}^{(2)\rho}$ $=$ $h^{\rho\sigma}(\partial_{\mu}\tau_{\sigma}{}^{a}-\partial_{\sigma}\tau_{(\mu}{}^{a})\tau_{\nu)}{}^{b}\eta_{ab}\,$, (3.2)
+[…]
+$\Gamma_{\mu\nu}^{(-2)\rho}$ $=$ $-\partial_{(\mu}\tau^{\sigma a}h_{\nu)\sigma}\tau^{\rho}{}_{a}-\tau^{\rho a}\tau^{\sigma}{}_{a}\partial_{\sigma}h_{\mu\nu}\,.$ (3.4)
+Inspecting the transformation rules under infinitesimal diffeomorphisms given by the Lie derivative $L_{\xi}$, i.e.,
+$\delta_{\xi}h_{\mu\nu}$ $=$ $L_{\xi}h_{\mu\nu}=\xi^{\rho}\partial_{\rho}h_{\mu\nu}+2\partial_{(\mu}\xi^{\rho}h_{\nu)\rho}\,$, (3.5)
+$\delta_{\xi}h^{\mu\nu}$ $=$ $L_{\xi}h^{\mu\nu}=\xi^{\rho}\partial_{\rho}h^{\mu\nu}-2\partial_{\rho}\xi^{(\mu}h^{\nu)\rho}\,$, (3.6)
+$\delta_{\xi}\tau_{\mu}{}^{a}$ $=$ $L_{\xi}\tau_{\mu}{}^{a}=\xi^{\rho}\partial_{\rho}\tau_{\mu}{}^{a}+\partial_{\mu}\xi^{\rho}\tau_{\rho}{}^{a}\,$, (3.7)
+$\delta_{\xi}\tau^{\mu}{}_{a}$ $=$ $L_{\xi}\tau^{\mu}{}_{a}=\xi^{\rho}\partial_{\rho}\tau^{\mu}{}_{a}-\partial_{\rho}\xi^{\mu}\tau^{\rho}{}_{a}\,$, (3.8)
+where $\xi^{\mu}$ is an arbitrary parameter.
+It is straightforward to check that $\Gamma_{\mu\nu}^{(2)\rho}$ and $\Gamma_{\mu\nu}^{(-2)\rho}$ transform covariantly under infinitesimal diffeomorphisms and they can be interpret as covariant shifts of the non-covariant part of the non-relativistic connection, given by $\Gamma_{\mu\nu}^{(0)\rho}$.
 3.2 Covariant derivative
@@
 $\nabla_{\mu}\tau_{\nu}{}^{a}$ $=$ $\partial_{\mu}\tau_{\nu}{}^{a}-\Gamma_{\mu\nu}^{\rho}\tau_{\rho}{}^{a}$
-$=$ $\partial_{\mu}\tau_{\nu}{}^{a}+\partial_{(\mu}{\tau_{\nu)}{}^{a}}-\frac{1}{c^{2}}\partial_{(\mu}{\tau_{\nu)}{}^{a}}+T_{\mu\nu}^{a}\,,$
-where $T_{\mu\nu}{}^{a}$ is given by the following covariant quantity,
-$T_{\mu\nu}{}^{a}$ $=$ $\eta_{bc}\eta^{ad}\nabla_{(\mu|}{\tau_{\rho}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}-\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu|}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}$ (3.2)
-$-\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{(\mu}{\tau_{\rho}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}+\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}\,.$
-An interesting observation of $\nabla_{\mu}\tau_{\nu}{}^{a}$ is that this quantity does not diverge, but contains $\frac{1}{c^{2}}$ that we need to track in order to correctly write the finite Lagrangian.
-This means that the finite Lagrangian will evolve terms with powers of $c$, but this is just the price to pay for writing it in covariant form.
-The antisymmetric part of $\partial_{\mu}\tau_{\nu}^{a}$ satisfies
-$\nabla_{[\mu}\tau_{\nu]}^{a}$ $=$ $\partial_{[\mu}\tau_{\nu]}{}^{a}$ (3.3)
-while the symmetric part can be written as
-$\partial_{(\mu}\tau_{\nu)}{}^{a}=\frac{c^{2}}{2c^{2}-1}\Big{(}\nabla_{(\mu}\tau_{\nu)}{}^{a}-T_{\mu\nu}{}^{a}\Big{)}=Q_{\mu\nu}{}^{a}\,.$ (3.4)
-On the other hand the covariant derivative of $\tau^{\nu}{}_{a}$ is given by
-$\nabla_{\mu}\tau^{\nu}{}_{a}=\partial_{\mu}\tau^{\nu}{}_{a}+\Gamma_{\mu\rho}^{\nu}\tau^{\rho}{}_{a}=\frac{1}{2}\partial_{\mu}\tau^{\nu}{}_{a}+S_{\mu}{}^{\nu}{}_{a}\,,$ (3.5)
-where $S_{\mu}{}^{\nu}{}_{a}$ is a covariant quantity given by
-$S_{\mu}{}^{\nu}{}_{a}$ $=$ $c^{2}\Big{(}\nabla_{[\rho}{\tau_{\sigma]}^{b}}\eta_{bc}\tau_{\mu}^{c}\tau^{\rho}_{a}h^{\nu\sigma}+\nabla_{[\mu}{\tau_{\rho]}^{b}}\eta_{ab}h^{\nu\rho}\Big{)}+\frac{1}{2}Q_{\mu\rho}{}^{b}(\tau^{\nu}_{b}\tau^{\rho}_{a}+\eta_{ab}h^{\nu\rho})$ (3.6)
-$+\frac{1}{2}\nabla_{[\mu}{\tau_{\rho]}^{b}}(2\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d}-\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}h^{\nu\rho})+\nabla_{[\rho}{\tau_{\sigma]}^{b}}(\eta_{bc}\tau_{\mu}^{c}\tau^{\sigma}_{a}h^{\nu\rho}+\eta_{bc}\eta^{de}\tau_{\mu}^{c}\tau^{\nu}_{d}\tau^{\rho}_{a}\tau^{\sigma}_{e})$
-$+\frac{1}{4c^{2}}\Big{(}Q_{\mu\rho}{}^{b}(-2\tau^{\nu}_{b}\tau^{\rho}_{a}+2\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})+\nabla_{\mu}{\tau_{\rho}^{b}}(\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})$
-$-\nabla_{\rho}{\tau_{\mu}^{b}}(\tau^{\nu}_{b}\tau^{\rho}_{a}-\eta_{ab}\eta^{cd}\tau^{\nu}_{c}\tau^{\rho}_{d})+4\nabla_{[\rho}{\tau_{\sigma]}^{b}}\eta_{bc}\eta^{de}\tau_{\mu}^{c}\tau^{\nu}_{d}\tau^{\sigma}_{a}\tau^{\rho}_{e}\Big{)}\,.$
-Considering the Newton-Cartan relation, it is easy to prove that the ungauged Lagrangian can be written in terms of $\nabla_{\mu}\tau_{\nu}{}^{a}$, $\nabla_{\mu}\tau^{\nu}{}_{a}$ and the tensors $Q_{\mu\nu}{}^{a}$ and $S_{\mu}{}^{\nu}{}_{a}$, which also depends on the previous covariant derivatives.
-In other words, we do not need to compute the covariant derivative of $h_{\mu\nu}$ and $h^{\mu\nu}$ to write, for example, the Riemann tensor, the Ricci tensor or the Ricci scalar in covariant form.
+$=$ $\partial_{\mu}\tau_{\nu}{}^{a}+T_{\mu\nu}^{a}\,$,
+where $T_{\mu\nu}{}^{a}=T_{\nu\mu}{}^{a}$ is given by the following quantity,
+$T_{\mu\nu}{}^{a}$ $=$ $+\partial_{(\mu}{\tau_{\nu)}{}^{a}}+\eta_{bc}\eta^{ad}\nabla_{(\mu|}{\tau_{\rho}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}-\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu|}{}^{b}}\tau_{|\nu)}{}^{c}\tau^{\rho}{}_{d}$ (3.10)
+$-\frac{1}{c^{2}}\partial_{(\mu}{\tau_{\nu)}{}^{a}}-\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{(\mu}{\tau_{\rho}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}+\frac{1}{c^{2}}\eta_{bc}\eta^{ad}\nabla_{\rho}{\tau_{(\mu}{}^{b}}\tau_{\nu)}{}^{c}\tau^{\rho}{}_{d}\,$.
+An interesting observation of $\nabla_{\mu}\tau_{\nu}{}^{a}$ is that this quantity does not diverge, but contains $\frac{1}{c^{2}}$ terms that, in general, one needs to track in order to correctly write a finite Lagrangian.
+Moreover, the antisymmetric part of $\partial_{\mu}\tau_{\nu}^{a}$ satisfies
+$\nabla_{[\mu}\tau_{\nu]}^{a}$ $=$ $\partial_{[\mu}\tau_{\nu]}{}^{a}\,$. (3.11)
+Therefore, the antisymmetric part of $\partial_{\mu}\tau_{\nu}^{a}$ is covariant under diffeomorphims, giving rise to the well-known intrinsic torsion of the theory [13].
+In order to write the two-derivative ungauged action in a covariant way we will just need $R^{(0)}$ given in (LABEL:Scalar) (which we prove in the appendix that it is a covariant quantity), and the relation (3.11) to replace ordinary derivatives by covariant derivatives in some antisymmetric expressions in the terms coming from the $\bar{H}^{2}$ contribution.
 3.3 Gauge transformations
-Transformation rules under the non-Abelian gauge symmetry can be extracted from (1.9), (1.10) and (1.11).
+The transformation rules under the non-Abelian gauge symmetry can be extracted from (1.13) and (1.14).
 We will impose an extra expansion on the gauge parameter and the structure constants following [22],
@@
 At this point, one can think that the rescaling of the structure constants might include divergences in the gauge Lagrangian.
-However, this reescaling is safe because higher order $\hat{f}_{ijk}$ contributions are $\propto\frac{1}{c}$.
+However, this rescaling is safe because higher order $\hat{f}_{ijk}$ contributions are $\propto\frac{1}{c}$.
 Then, the rescaling only introduces cubic contributions in gauge fields in the finite Lagrangian, which we will present in the next section.
-After exploring the gauge transformation (1.9), (1.10) and (1.11) we find the following rules upon taking the limit $c\rightarrow\infty$,
+After exploring the gauge transformations (1.13) and (1.14) we find the following rules when taking the limit $c\rightarrow\infty$,
 $\delta_{\lambda}a_{\mu}^{i}$ $=$ $\partial_{\mu}\lambda^{i}+f^{i}{}_{jk}\lambda^{j}a_{\mu}^{k}\,$, (3.14)
 $\delta_{\lambda}\alpha_{-}^{i}$ $=$ $f^{i}{}_{jk}\lambda^{j}\alpha_{-}^{k}\,$, (3.15)
-$\delta_{\lambda}b_{\mu\nu}$ $=$ $-(\partial_{[\mu}\lambda^{i})\alpha_{-i}\tau_{\nu]}^{-}\,,$ (3.11)
-$\delta_{\lambda}h_{\mu\nu}$ $=$ $0\,,$ (3.12)
-$\delta_{\lambda}h^{\mu\nu}$ $=$ $0\,,$ (3.13)
-$\delta_{\lambda}\tau_{\mu}{}^{a}$ $=$ $0\,,$ (3.14)
-$\delta_{\lambda}\tau^{\mu}{}_{a}$ $=$ $0\,.$ (3.15)
+$\delta_{\lambda}b_{\mu\nu}$ $=$ $-(\partial_{[\mu}\lambda^{i})\alpha_{-i}\tau_{\nu]}^{-}\,,\,$, (3.16)
+while the remaining fields do not transform under gauge transformations.
 From the previous transformations, we learn that $a_{\mu}^{i}$ transforms as a gauge connection, $\alpha_{-}^{i}$ transforms as a vector and $b_{\mu\nu}$ transforms in a non-covariant way, emulating a Green-Schwarz mechanism.
-Interestingly, this mechanism cannot be eliminated through field redefinitions in the relativistic scenario.
-However, when one considers the NR limit, it is possible to define the field
-$\bar{b}_{\mu\nu}=b_{\mu\nu}+a_{[\mu|i}\alpha_{-}^{i}\tau_{|\nu]-}$ (3.16)
+While in the relativistic case this transformation is non-ambiguous and cannot be removed, when one considers the NR expansions presented in this work (based on [22]), it is possible to define the field
+$\tilde{b}_{\mu\nu}=b_{\mu\nu}+a_{[\mu|i}\alpha_{-}^{i}\tau_{|\nu]-}$ (3.17)
 which is gauge invariant, i.e.
-$\delta_{\lambda}\bar{b}_{\mu\nu}=0$.
-This is an interesting property of the heterotic supergravity under the NR limit which will help us to construct a suitable curvature for the $\bar{b}$-field.
+$\delta_{\lambda}\tilde{b}_{\mu\nu}=0$.
+This is an interesting property of the heterotic supergravity under the NR limit, which will help us to construct a suitable curvature for the $\tilde{b}-$field and, also, is potentially telling us that the cancellation of anomalies could be trivial under this limit.
+If this is the case, one could construct different formulations of heterotic supergravity using arbitrary gauge groups, beyond $SO(32)$ and $E8\times E8$.
+The NR NS-NS supergravity is invariant under boost transformations given by [13],
+$\delta_{\lambda}h_{\mu\nu}$ $=$ $-2\lambda_{aa^{\prime}}e_{(\mu}{}^{a^{\prime}}\tau_{\nu)}{}^{a}\,$, (3.18)
+$\delta_{\lambda}\tau^{\mu\nu}$ $=$ $2\lambda_{a}{}^{a^{\prime}}e^{(\mu}{}_{a^{\prime}}\tau^{\nu)a}\,$, (3.19)
+$\delta_{\lambda}b_{\mu\nu}$ $=$ $-2\epsilon_{ab}\lambda^{a}{}_{a^{\prime}}\tau_{[\mu}{}^{b}e_{\nu]}{}^{a^{\prime}}\,$. (3.20)
+Since the gauge fields do not transform under this symmetry, the field redefinition (3.17) do not deform the boost transformation of the new $\tilde{b}_{\mu\nu}$ field.
 3.4 Gauging the covariant derivative
@@
 $f_{\mu\nu}{}^{i}=2\partial_{[\mu}a_{\nu]}{}^{i}-f^{i}{}_{jk}a_{\mu}{}^{j}a_{\nu}{}^{k}\,$, (3.22)
-The field $\alpha_{-}^{i}$ is already covariant, so we do not need to construct its curvature. If we consider $\bar{b}$ as the fundamental field of the theory, then we define its curvature as
-$\bar{h}_{\mu\nu\rho}=3\partial_{[\mu}\bar{b}_{\nu\rho]}\,=h_{\mu\nu\rho}-3c^{(g)}_{\mu\nu\rho}.$ (3.19)
+The field $\alpha_{-}^{i}$ is already covariant, so we do not need to construct its curvature. If we consider $\tilde{b}$ as the fundamental field of the theory, then we define its curvature as
+$\bar{h}_{\mu\nu\rho}=3\partial_{[\mu}\tilde{b}_{\nu\rho]}\,=h_{\mu\nu\rho}-3c^{(g)}_{\mu\nu\rho}$. (3.23)
 From the previous curvature, we can identify the equivalent quantity to the Chern-Simons terms in this limit,
 $c^{(g)}_{\mu\nu\rho}=-\partial_{[\mu}(a_{\nu}^{i}\tau_{\rho]}^{-}\alpha_{-i}).$ (3.24)
-In the next section, we will use these curvatures to present the finite NR gauge Lagrangian.
+The introduction of the $\tilde{b}_{\mu\nu}$ field is a crucial result of this paper.
+It shows that the gauge Green-Schwarz mechanism is trivial in the NR-limit of the heterotic supergravity.
+Furthermore, the Chern-Simons terms in form language are given by the exterior derivative of a two form, simplifying the Bianchi identity when we compare to the relativistic case.
+In section (4), we will use these curvatures to present the finite NR gauge Lagrangian.
 4 Finite Lagrangian
-In this section we construct all the finite contributions to the heterotic supergravity Lagrangian in covariant form upon taking the NR limit.
-For the readers interested only in the final result, we report the final form of the gauge Lagrangian in the conclusions.
+In this section, we construct all the finite contributions to the heterotic supergravity Lagrangian in covariant form upon taking the NR limit.
+For readers interested only in the final result, we report the final form of the gauge Lagrangian at the end of this section.
 4.1 Ungauged Lagrangian
-We would like to present the finite heterotic Lagrangian in terms of covariant quantities.
-We will choose objects which are covariant under infinitesimal diffeomorphisms.
-Due to the previous section, we have all the pieces to write the ungauged part of the finite Lagrangian in its covariant form.
-This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term and the H-term given by
-$L$ $=$ $R^{(0)}+4\partial_{\mu}\varphi\partial^{\mu}\varphi-\frac{1}{12}H^{(0)}_{\mu\nu\rho}H^{(0)}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$ (4.1)
-$-\frac{1}{2}H^{(0)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}H^{(2)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\eta^{cd}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c}\tau^{\sigma}{}_{b}\tau^{\gamma}{}_{d}h^{\rho\epsilon}\,,$
+This Lagrangian contains contributions coming from the Ricci scalar, the dilaton term, and the H-term, given by
+$L$ $=$ $ee^{-2\varphi}(R^{(0)}+4\partial_{\mu}\varphi\partial^{\mu}\varphi-\frac{1}{12}H^{(0)}_{\mu\nu\rho}H^{(0)}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$ (4.1)
+$-\frac{1}{2}H^{(0)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{4}H^{(2)}_{\mu\nu\rho}H^{(2)}_{\sigma\gamma\epsilon}\eta^{ab}\eta^{cd}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c}\tau^{\sigma}{}_{b}\tau^{\gamma}{}_{d}h^{\rho\epsilon})\,$,
 where
 $H^{(0)}_{\mu\nu\rho}$ $=$ $h_{\mu\nu\rho}=3\partial_{[\mu}b_{\nu\rho]}\,$, (4.2)
-$H^{(2)}_{\mu\nu\rho}$ $=$ $-6c^{2}\epsilon_{ab}\nabla_{[\mu}\tau_{\nu}{}^{a}\tau_{\rho]}{}^{b}\,,$ (4.3)
-and the Ricci scalar is given by
-$R^{(0)}$ $=$ $8\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau^{\mu}{}_{a}}h^{\nu\rho}+5\nabla_{[\mu}{\tau_{\nu]}{}^{c}}\eta^{ab}(\nabla_{\rho}{\tau^{\mu}{}_{a}}\tau^{\nu}{}_{c}\tau^{\rho}{}_{b}+\nabla_{\rho}{\tau^{\rho}{}_{a}}\tau^{\mu}{}_{b}\tau^{\nu}{}_{c})$ (4.4)
-$-\frac{3}{8}\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}(\tau^{\nu}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\rho}-2\tau^{\mu}{}_{a}\tau^{\sigma}{}_{b}h^{\nu\rho}+\tau^{\mu}{}_{a}\tau^{\rho}{}_{b}h^{\nu\sigma})$
-$+\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}h^{\mu\rho}h^{\nu\sigma}+4c^{2}\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}h^{\mu\sigma}h^{\nu\rho}$
-$+\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}\eta^{cd}(2\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}-\tau^{\mu}{}_{c}\tau^{\rho}{}_{d}h^{\nu\sigma}-\tau^{\nu}{}_{c}\tau^{\sigma}{}_{d}h^{\mu\rho})$
-$+\frac{1}{c^{2}}\nabla_{\mu}{\tau_{\nu}{}^{a}}\nabla_{\rho}{\tau_{\sigma}{}^{b}}\eta_{ab}\eta^{cd}(\tau^{\nu}{}_{c}\tau^{\sigma}{}_{d}h^{\mu\rho}-2\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}+\tau^{\mu}{}_{c}\tau^{\rho}{}_{d}h^{\nu\sigma})$
-$-2\nabla_{\mu}{\nabla_{\nu}{\tau_{\rho}{}^{a}}}\tau^{[\rho}{}_{a}h^{\nu]\mu}+f(Q,S)\,,$
-where $f(Q,S)$ is a covariant contribution since both $Q_{\mu\nu}{}^{a}$ and $S_{\mu}{}^{\nu}{}_{a}$ are covariant,
-$f(Q,S)$ $=$ $8Q_{\mu\nu}{}^{a}(S_{\rho}{}^{[\mu}{}_{a}h^{\rho]\nu}+\nabla_{\rho}{\tau^{[\rho}{}_{a}}h^{\mu]\nu})-8\nabla_{[\nu}{\tau_{\rho]}{}^{a}}S_{\mu}{}^{\nu}{}_{a}h^{\mu\rho}-\frac{1}{2}Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\tau^{\mu}{}_{a}\tau^{\rho}{}_{b}h^{\nu\sigma}$ (4.5)
-$+Q_{\mu\nu}{}^{a}\eta^{bc}(-2\nabla_{\rho}{\tau^{\mu}{}_{b}}\tau^{\nu}{}_{c}\tau^{\rho}{}_{a}+\nabla_{\rho}{\tau^{\mu}{}_{b}}\tau^{\nu}{}_{a}\tau^{\rho}{}_{c}+\nabla_{\rho}{\tau^{\rho}{}_{b}}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c})$
-$+Q_{\mu\nu}{}^{a}\eta^{bc}(2S_{\rho}{}^{\mu}{}_{b}\tau^{\nu}{}_{c}\tau^{\rho}{}_{a}-S_{\rho}{}^{\mu}{}_{b}\tau^{\nu}{}_{a}\tau^{\rho}{}_{c}-S_{\rho}{}^{\rho}{}_{b}\tau^{\mu}{}_{a}\tau^{\nu}{}_{c})$
-$+\nabla_{\rho}{\tau_{\sigma}{}^{b}}Q_{\mu\nu}{}^{a}(-\frac{3}{2}\tau^{\rho}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\nu}+2\tau^{\mu}{}_{a}\tau^{[\rho}{}_{b}h^{\sigma]\nu}+2\tau^{\rho}{}_{b}\tau^{\sigma}{}_{a}h^{\mu\nu}+4\tau^{\mu}{}_{b}\tau^{[\rho}{}_{a}h^{\sigma]\nu})$
-$+\nabla_{\nu}{\tau_{\rho}{}^{c}}\eta^{ab}(-\frac{5}{2}S_{\mu}{}^{\nu}{}_{a}\tau^{\mu}{}_{b}\tau^{\rho}{}_{c}+5S_{\mu}{}^{\mu}{}_{a}\tau^{[\nu}{}_{c}\tau^{\rho]}{}_{b})$
-$+\frac{5}{2}\nabla_{\rho}{\tau_{\nu}{}^{c}}\eta^{ab}S_{\mu}{}^{\nu}{}_{a}\tau^{\mu}{}_{b}\tau^{\rho}{}_{c}+2Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\eta_{ab}h^{\mu[\nu}h^{\rho]\sigma}+2Q_{\mu\nu}{}^{a}Q_{\rho\sigma}{}^{b}\eta_{ab}\eta^{cd}\tau^{\mu}{}_{c}\tau^{[\rho}{}_{d}h^{\nu]\sigma}$
-$+3\nabla_{\rho}{\tau_{\sigma}{}^{b}}Q_{\mu\nu}{}^{a}\eta_{ab}\eta^{cd}\tau^{\mu}{}_{c}\tau^{[\sigma}{}_{d}h^{\rho]\nu})+\frac{4}{c^{2}}\nabla_{[\rho}{\tau_{\sigma]}{}^{b}}Q_{\mu\nu}{}^{a}(\tau^{\rho}{}_{a}\tau^{\sigma}{}_{b}h^{\mu\nu}+\tau^{\mu}{}_{b}\tau^{\sigma}{}_{a}h^{\nu\rho})$
-$+4\nabla_{\mu}{Q_{\nu\rho}{}^{a}}\tau^{[\mu}{}_{a}h^{\nu]\rho}\,.$
-In $R^{(0)}$ we can explicitly see that terms proportional to $c^{2}$ and $\frac{1}{c^{2}}$ naturally arises when we write this object using covariant derivatives.
-When we replace the covariant derivatives by their definitions, all these divergent and convergent quantities cancel, giving a finite contribution in terms of partial derivatives.
+$H^{(2)}_{\mu\nu\rho}$ $=$ $-6c^{2}\epsilon_{ab}\nabla_{[\mu}\tau_{\nu}{}^{a}\tau_{\rho]}{}^{b}\,$. (4.3)
+Collecting all the contributions coming from the H-terms, we find the following finite Lagrangian
+$L=ee^{-2\varphi}(R^{(0)}+4\partial_{\mu}\varphi\partial_{\nu}\varphi h^{\mu\nu}+T_{H^{2}})$ (4.4)
+where
+$T_{H^{2}}$ $=$ $2\eta_{ab}\eta^{cd}\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{[\rho}{\tau_{\sigma]}{}^{b}}\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}+2\epsilon_{ab}\epsilon_{cd}\eta^{ae}\eta^{cf}\nabla_{[\mu}{\tau_{\nu]}{}^{b}}\nabla_{[\rho}{\tau_{\sigma]}{}^{d}}\tau^{\nu}{}_{f}\tau^{\sigma}{}_{e}h^{\mu\rho}$ (4.5)
+$-\frac{1}{12}h_{\mu\nu\rho}h_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\epsilon_{ab}\eta^{ac}\nabla_{[\mu}{\tau_{\nu]}{}^{b}}h_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\rho}h^{\nu\gamma}$
+The Ricci scalar can be easily computed considering the non-relativistic expansions (1.1)-(1.2). The generic form of this object is as follows,
+$R=c^{2}R^{(2)}+R^{(0)}+\frac{1}{c^{2}}R^{-2}+\frac{1}{c^{4}}R^{-4}\,.$ (4.6)
+Every contribution in the previous expansion transforms as a scalar, as the reader can easily check (we attach the explicit demostration for $R^{(0)}$ in the appendix C).
+The Lagrangian (4.4), together with (4.5) and $R^{(0)}$, constitutes the ungauged bosonic part of the heterotic supergravity.
 4.2 Gauged terms
-Before analyzing the gauge contributions of the heteorotic Lagrangian, let’s focus on the ungauged terms depending on the b-field,
-$L_{b}$ $=$ $-\frac{1}{4}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{b_{\gamma\epsilon}}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\sigma}{}_{c}h^{\mu\rho}h^{\nu\gamma}+\frac{1}{2}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{b_{\gamma\epsilon}}h^{\mu\gamma}h^{\nu\sigma}h^{\rho\epsilon}$ (4.6)
-$-\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\rho}{}_{c}h^{\mu\sigma}h^{\nu\gamma}-\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}{}^{b}}\partial_{\rho}{b_{\sigma\gamma}}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}\,.$
-These terms are not gauge invariant, due to the transformation (3.11). Therefore, we rewrite these contributions in terms of the new curvature for the b-field,
-$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}{}^{b}\bar{h}_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}+\nabla_{\mu}\tau_{\nu}^{-}f_{\gamma\rho}{}^{i}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}+L_{\textrm{comp}}\,,$ (4.7)
-where $L_{\textrm{comp}}=L^{(2)}_{\textrm{non-cov}}+L^{(3)}_{\textrm{non-cov}}+L^{(4)}_{\textrm{non-cov}}$ with
-$L^{(2)}_{\textrm{non-cov}}$ $=$ $\nabla_{\mu}\tau_{\nu}^{-}a_{[\rho}^{i}\nabla_{\gamma]}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}+3\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}a_{[\sigma}^{i}\alpha_{-i}\nabla_{\rho}\tau_{\gamma]}^{-}\tau^{\sigma}_{c}h^{\mu\gamma}h^{\nu\rho}$ (4.8)
-$+\frac{1}{2}h_{\mu\nu\rho}a_{\gamma}^{i}\alpha_{-i}\nabla_{\sigma}(\tau_{\epsilon}^{-})h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}\,,$
-$L^{(3)}_{\textrm{non-cov}}=-\nabla_{\mu}\tau_{\nu}^{-}f^{i}{}_{jk}a_{\gamma}^{j}a_{\rho}{}^{k}\alpha_{-i}h^{\mu\gamma}h^{\nu\rho}\,,$ (4.9)
-and
-$L^{(4)}_{\textrm{non-cov}}$ $=$ $\frac{3}{4}a_{[\nu}^{i}\alpha_{-i}\nabla_{\mu}(\tau_{\rho]^{-}})a_{\gamma}^{j}\alpha_{-j}\nabla_{\sigma}\tau_{\epsilon}^{-}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}\,.$ (4.10)
-In the next subsection we will study the contributions to the gauge Lagrangian coming from the Yang-Mills terms and the Chern-Simons contributions.
-We split the presentation according to an expansion in gauge fields.
-4.3 Quadratic gauged contributions
-4.3.1 Quadratic Yang-Mills contributions
-The quadratic contributions coming from $-\frac{1}{4}\hat{F}^{2}$ are given by
-$-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$
-$+2\alpha_{-i}\eta^{ab}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau_{\sigma}^{-}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-2\alpha_{-i}\nabla_{\mu}\tau_{\nu}{}^{-}\nabla_{[\rho}a_{\sigma]}^{i}h^{\mu\rho}h^{\nu\sigma}\,.$ (4.11)
-The previous expression can be written in a covariant form as
-$-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$
-$-2\alpha_{-i}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau^{\nu}_{+}h^{\mu\rho}-\alpha_{-i}\nabla_{\mu}\tau_{\nu}{}^{-}f_{\rho\sigma}^{i}h^{\mu\rho}h^{\nu\sigma}+L^{(3)}(a,\alpha)$ (4.12)
-where the inclusion of the (gauge) covariant derivative and the curvature $f_{\mu\nu i}$ introduces cubic contributions
-$L^{(3)}(a,\alpha)=-\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}f^{i}{}_{jk}a^{j}_{\rho}a^{k}_{\sigma}h^{\mu\rho}h^{\nu\sigma}+2\alpha_{-i}\eta^{ab}f^{i}{}_{jk}a_{\mu}^{j}\alpha_{-}^{k}\nabla_{[\nu}\tau_{\rho]}^{-}\tau_{\sigma}^{-}\tau_{a}^{\nu}\tau_{b}^{\sigma}h^{\mu\rho}\,.$ (4.13)
-4.3.2 Quadratic Chern-Simons contributions
-The quadratic contributions coming from $-\frac{1}{12}\bar{H}^{2}$ are given by
-$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}$
-$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}$
-$-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}a_{\sigma]i}h^{\mu\sigma}h^{\nu\rho}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}h_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}-L^{(2)}_{\textrm{non-cov}}\,.$ (4.14)
-Among these terms we have identified all the quadratic non-covariant contributions needed in (4.7). We can write the previous expression in covariant form as
-$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}$
-$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}$
-$-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}f_{\rho\sigma i}h^{\mu\sigma}h^{\nu\rho}-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}\bar{h}_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}-L^{(2)}_{\textrm{non-cov}}+\tilde{L}^{(3,4)}(a,\alpha)\,,$ (4.15)
-where once again the inclusion of the curvatures $f_{\mu\nu i}$ and $\bar{h}_{\rho\sigma\gamma}$ in the last line of the previous expression introduces cubic and quartic contributions, respectively, that we will take into account when we analyze the next orders.
-We capture all these new contributions under the notation $\tilde{L}^{(3,4)}(a,\alpha)$, which splits in the following form
-$\tilde{L}^{(3)}(a,\alpha)=-\alpha_{-}^{i}\nabla_{\mu}\tau_{\nu}^{-}f_{i}{}^{jk}a_{\rho j}a_{\sigma k}h^{\mu\sigma}h^{\nu\rho}\,,$ (4.16)
-and
-$\tilde{L}^{(4)}(a,\alpha)=\frac{3}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}a_{[\sigma}^{j}\alpha_{-j}\nabla_{\rho}(\tau_{\gamma]}^{-})\tau_{+}^{\rho}h^{\mu\sigma}h^{\nu\gamma}\,.$ (4.17)
-4.4 Cubic gauge contributions
-4.4.1 Cubic Yang-Mills contributions
-The cubic contributions coming from $-\frac{1}{4}\hat{F}^{2}$ are given by
-$-L^{(3)}_{\textrm{non-cov}}-L^{(3)}(a,\alpha)-\tilde{L}^{(3)}(a,\alpha)\,.$ (4.18)
-The previous expression exactly cancels the contributions obtained in (4.13) and (4.16) and there are no additional contributions at this order in fields.
-4.5 Quartic gauge contributions
-4.5.1 Quartic Chern-Simons contributions
-The quartic terms in $-\frac{1}{12}\bar{H}^{2}$ are given by
-$-\tilde{L}^{(4)}(a,\alpha)\,,$ (4.19)
-and there are no other additional contributions at this order, or higher, in fields.
+Now we compute the gauge contributions of the heterotic Lagrangian given by the inclusion of the $\hat{A}_{\mu}{}^{i}$ field.
+[…]
+with
+$\bar{H}_{\mu\nu\rho}=3\left(\partial_{[\mu}\hat{B}_{\nu\rho]}-\hat{C}_{\mu\nu\rho}^{(g)}\right)\,$, (4.8)
+$\hat{C}_{\mu\nu\rho}^{(g)}=\hat{A}^{i}_{[\mu}\partial_{\nu}\hat{A}_{\rho]i}-\frac{1}{3}\hat{f}_{ijk}\hat{A}_{\mu}^{i}\hat{A}_{\nu}^{j}\hat{A}_{\rho}^{k}\,$, (4.9)
+$\hat{F}_{\mu\nu}{}^{i}=2\partial_{[\mu}\hat{A}_{\nu]}{}^{i}-\hat{f}^{i}{}_{jk}\hat{A}_{\mu}{}^{j}\hat{A}_{\nu}{}^{k}\,$, (4.10)
+and $\hat{f}_{ijk}=cf_{ijk}$. In this case, the Lagrangian also remains finite after considering the limit $c\rightarrow\infty$ and it can be written as
+$L=ee^{-2\varphi}\Big(R^{(0)}+4\partial_{\mu}\varphi\partial_{\nu}\varphi h^{\mu\nu}+T_{H^{2}}+T_{F^{2}}\Big)$ (4.11)
+where
+$T_{H^{2}}=\frac{1}{2}\alpha_{-}^{2}\eta^{ab}\nabla_{[\mu}{\tau_{\nu]}^{-}}\bar{h}_{\rho\sigma\gamma}\tau_{\epsilon}^{-}\tau^{\rho}{}_{a}\tau^{\epsilon}{}_{b}h^{\mu\sigma}h^{\nu\gamma}+2\alpha_{-}^{2}\epsilon_{ab}\eta^{ac}\eta^{de}\nabla_{[\mu}{\tau_{\nu]}^{b}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau_{\gamma}^{-}\tau^{\mu}_{d}\tau^{\sigma}_{c}\tau^{\gamma}_{e}h^{\nu\rho}$
+$-\frac{1}{2}\alpha_{-}^{4}\eta^{ab}\eta^{cd}\nabla_{[\mu}{\tau_{\nu]}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau_{\gamma}^{-}\tau_{\epsilon}^{-}\tau^{\mu}_{a}\tau^{\sigma}_{c}\tau^{\gamma}_{d}\tau^{\epsilon}_{b}h^{\nu\rho}$
+$+2\eta_{ab}\eta^{cd}\nabla_{[\mu}{\tau_{\nu]}{}^{a}}\nabla_{[\rho}{\tau_{\sigma]}{}^{b}}\tau^{\mu}{}_{c}\tau^{\sigma}{}_{d}h^{\nu\rho}+2\epsilon_{ab}\epsilon_{cd}\eta^{ae}\eta^{cf}\nabla_{[\mu}{\tau_{\nu]}{}^{b}}\nabla_{[\rho}{\tau_{\sigma]}{}^{d}}\tau^{\nu}{}_{f}\tau^{\sigma}{}_{e}h^{\mu\rho}$
+$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}{}^{b}}\tau^{\rho}{}_{c}\bar{h}_{\rho\sigma\gamma}h^{\mu\sigma}h^{\nu\gamma}$ (4.12)
+is the finite contribution coming from the $\bar{H}^{2}$-term, with $\alpha_{-}^{2}=\alpha_{-}^{i}\alpha_{-}^{j}\eta_{ij}$, and the gauge-invariant 3-form is given by
+$\bar{h}_{\mu\nu\rho}=h_{\mu\nu\rho}+3\partial_{[\mu}(a_{\nu}^{i}\tau_{\rho]}^{-}\alpha_{-i})$ (4.13)
+The finite contribution coming from the $\hat{F}^{2}$-term is given by
+$T_{F^{2}}$ $=$ $-2\alpha_{-}^{2}\eta^{ab}\nabla_{[\mu}{\tau_{\nu]}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-2\alpha_{-}^{i}\eta^{ab}\eta_{ij}\nabla_{\mu}{\alpha_{-}^{j}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{\sigma}_{b}h^{\mu\nu}$ (4.14)
+$-f_{\mu\nu}^{i}\alpha_{-}^{j}\eta_{ij}\nabla_{\rho}{\tau_{\sigma}^{-}}h^{\mu\rho}h^{\nu\sigma}-\frac{1}{2}\eta^{ab}\eta_{ij}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{\nu}{\alpha_{-}^{j}}\tau_{\rho}^{-}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{\sigma}_{b}h^{\mu\nu}\,$,
+As the reader can easily verify, all the gauge contributions to the heterotic Lagrangian in the NR limit depend on the vector $\alpha_{-}{}^{i}$.
+4.3 Comparison to the Bergshoeff-Romano formulation
+As an alternative to the expansion for the gauge field discussed in this paper, it is possible to construct a finite formulation by considering the following expansion [21],
+$V_{\nu}{}^{i}=c^{2}\tilde{\tau}_{\nu}{}^{-}v_{-}{}^{i}+\tilde{\tau}_{\nu}{}^{+}v_{+}{}^{i}+\tilde{e}_{\nu}{}^{a^{\prime}}v_{a^{\prime}}{}^{i}\,$, (4.15)
+where $V_{\mu}{}^{i}$ plays the role of the $A_{\mu}{}^{i}$ vector. In this formalism, the c-expansion of the vielbein in this formulation includes gauge field contributions and it is given by
+$E_{\mu}{}^{-}$ $=$ $c\tilde{\tau}_{\mu}{}^{-},$ (4.16)
+$\hat{E}_{\mu}{}^{+}$ $=$ $-c^{3}\frac{v_{-}{}^{i}v_{-i}}{2}\tilde{\tau}_{\mu}{}^{-}+c\tilde{\tau}_{\mu}{}^{+}\,$, (4.17)
+$E_{\mu}{}^{a^{\prime}}$ $=$ $\tilde{e}_{\mu}{}^{a^{\prime}}\,$, (4.18)
+while the expansion for the B-field is given by
+$B_{\mu\nu}=c^{2}\tilde{\tau}_{\mu}{}^{a}{\tilde{\tau}_{\nu}}^{b}\epsilon_{ab}+2c^{2}\tilde{\tau}_{[\mu}{}^{-}\tilde{e}_{\nu]}{}^{a^{\prime}}v_{-}{}^{i}v_{a^{\prime}i}+\tilde{b}_{\mu\nu}\,.$ (4.19)
+The gauge transformations of the gauge degrees of freedom are given by
+$\delta_{\lambda}v_{+}{}^{i}$ $=$ $\partial_{+}\lambda^{i}+\sqrt{2}gf_{jk}{}^{i}\lambda^{j}v_{+}{}^{k}\,$, (4.20)
+$\delta_{\lambda}v_{-}{}^{i}$ $=$ $\sqrt{2}gf_{jk}{}^{i}\lambda^{j}v_{-}{}^{k}\,$, (4.21)
+$\delta_{\lambda}v_{a^{\prime}}{}^{i}$ $=$ $\partial_{a}\lambda^{i}+\sqrt{2}gf_{jk}{}^{i}\lambda^{j}v_{a^{\prime}}{}^{k}\,$. (4.22)
+We can read from the previous transformations that $v_{+}{}^{i}$ and $v_{a}{}^{i}$ play the role of connections, while $v_{-}{}^{i}$ is a gauge vector.
+In this prescription $\tilde{\tau}_{-}{}^{a}$ is gauge invariant but $\tilde{\tau}_{+}{}^{a}$ transforms in a non-covariant way under gauge transformations,
+$\delta\tilde{\tau}_{\mu}{}^{+}=\frac{v_{-}{}^{i}\partial_{-}\lambda^{i}}{1+v_{+}{}^{i}v_{-i}}\tilde{\tau}_{\mu}{}^{-}\,.$ (4.23)
+Since $v_{-}{}^{i}$ is boost-invariant and transforms as a vector under gauge transformations, we will work with the case $v_{-}{}^{i}=0$, so that all
+$\tilde{\tau}_{\mu}{}^{a}=\tau_{\mu}{}^{a}\quad\quad\tilde{e}_{\mu}{}^{a^{\prime}}=e_{\mu}{}^{a^{\prime}}\,$. (4.24)
+At this point, it is important to observe that taking $\alpha_{-}{}^{i}=0$ in the prescription used in the previous sections is not equivalent to $v_{-}{}^{i}=0$.
+In the formalism [21] this condition still produces a non-trivial Green-Schwarz mechanism for the b-field, given by
+$\delta_{\lambda}b_{\mu\nu}$ $=$ $2\tau_{[\mu}{}^{+}e_{\nu]}{}^{a^{\prime}}(v_{+}{}^{i}\partial_{a^{\prime}}\lambda_{i}-v_{a^{\prime}i}\partial_{+}\lambda^{i})-2e_{[\mu}{}^{a^{\prime}}e_{\nu]}{}^{b^{\prime}}v_{a^{\prime}}{}^{i}\partial_{b^{\prime}}\lambda_{i}$ (4.25)
+$-4(\tau_{[\mu}{}^{-}\tau_{\nu]}{}^{+}v_{+i}+\tau_{[\mu}{}^{-}e_{\nu]}{}^{a^{\prime}}v_{a^{\prime}i})\partial_{-}\lambda^{i}\,.$
+Moreover, in the case $v_{-}{}^{i}=0$, the transformation (4.25) is not ambiguous and it cannot be eliminated with a field redefinition.
+When $v_{-}{}^{i}\neq 0$, one can try to use the anomalous transformation of $\tilde{\tau}_{\mu}{}^{+}$ to trivialize the Green-Schwarz mechanism.
+The explicit finite gauged action in this formulation, for the case $v_{-}{}^{i}=0$ is written in the Appendix B.
 5 Discussion
-In this work we have studied the finite form of the heterotic supergravity Lagrangian, extending the formulation of [13].
-The Ricci scalar part of the Lagrangian has been rewritten in terms of $\nabla_{\mu}\tau_{\nu}{}^{a}$, $\nabla_{\mu}\tau^{\nu}{}_{a}$, $Q_{\mu\nu}{}^{a}$ (eq.
-(3.4)) and $S_{\mu}{}^{\nu}{}_{a}$ (eq.
-(3.6)), showing the explicit diffeomorphism invariance of the theory.
-When we include the heteoric gauge field, the bosonic part of the Lagrangian of this theory contains two divergent parts given by the Chern-Simons terms and the Yang-Mills Lagrangian which exactly compensate each other, as it happens in bosonic supergravity, giving rise to a finite Lagrangian.
-The gauge Lagrangian depends on a pair of gauge fields, $a_{\mu}^{i}$ and $\alpha_{-}^{i}$, the former transforming as a gauge connection, so we constructed its curvature, and the latter transforming as a gauge vector, so we do not need to construct a curvature.
-Interestingly enough, the Green-Schwarz mechanism for the b-field can be trivialized in this limit and therefore one can define $\tilde{b}_{\mu\nu}$ as in (3.16), and its curvature $\tilde{h}_{\mu\nu\rho}$, both of them gauge invariant.
-The gauge contributions in $\tilde{h}_{\mu\nu\rho}$ can be though of as Chern-Simons contributions.
-In this work we have presented the finite Lagrangian of the theory which contains quadratic, cubic and quartic contributions. The full gauge Lagrangian is given by
-$L^{(g)}_{\textrm{het}}$ $=$ $-\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\nu}_{a}\tau^{\sigma}_{b}h^{\mu\rho}-\alpha_{-i}\alpha_{-}^{i}\eta^{ab}\nabla_{\mu}{\tau_{\nu}^{-}}\nabla_{[\rho}{\tau_{\sigma]}^{-}}\tau^{\mu}_{a}\tau^{\rho}_{b}h^{\nu\sigma}$ (5.1)
-$-2\alpha_{-i}\nabla_{\mu}{\alpha_{-}^{i}}\nabla_{[\nu}{\tau_{\rho]}^{-}}\tau^{\nu}_{+}h^{\mu\rho}-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}^{b}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\mu}_{+}\tau^{[\sigma}_{c}h^{\rho]\nu}$
-$-\alpha_{-}^{i}\alpha_{-i}\epsilon_{ab}\eta^{ac}\nabla_{\mu}{\tau_{\nu}^{b}}\nabla_{\rho}{\tau_{\sigma}^{-}}\tau^{\nu}_{+}\tau^{[\rho}_{c}h^{\sigma]\mu}+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\sigma}_{b}\tau^{[\nu}_{a}h^{\mu]\rho}$
-$+\alpha_{-}^{i}\alpha_{-i}\eta^{ab}\nabla_{\mu}\tau_{\nu}^{-}\nabla_{\rho}\tau_{\sigma}^{-}\tau^{\rho}_{a}\tau^{[\mu}_{b}h^{\nu]\sigma}-\alpha_{-}^{i}\nabla_{\mu}{\tau_{\nu}^{-}}f_{\rho\sigma i}h^{\mu\sigma}h^{\nu\rho}$
-$-\frac{1}{2}\alpha_{-}^{i}\alpha_{-i}\nabla_{\mu}\tau_{\nu}^{-}\bar{h}_{\rho\sigma\gamma}\tau^{\rho}_{+}h^{\mu\sigma}h^{\nu\gamma}\,,$
-while the ungauged action must include the Chern-Simons terms in the curvature of the b-field as we show in (4.7), i.e.,
-$-\frac{1}{12}\bar{h}_{\mu\nu\rho}\bar{h}_{\sigma\gamma\epsilon}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\epsilon_{ab}\eta^{ac}\nabla_{\mu}\tau_{\nu}{}^{b}\bar{h}_{\rho\sigma\gamma}\tau^{\sigma}{}_{c}h^{\mu\gamma}h^{\nu\rho}\,.$ (5.2)
+In this work, we study the finite form of the heterotic supergravity Lagrangian, extending the formulation of [13], and providing an alternative way to that of [21] of taking this limit.
+The bosonic part of the Lagrangian of this theory contains two divergent parts given by the Chern-Simons terms and the Yang-Mills Lagrangian which exactly compensate each other, as it happens in bosonic supergravity, giving rise to a finite Lagrangian.
+The gauge Lagrangian depends on a pair of gauge fields, $a_{\mu}^{i}$ and $\alpha_{-}^{i}$, the former transforming as a gauge connection, so we construct its curvature, and the latter transforms as a gauge vector, so we do not need to construct a curvature.
+While the formulation presented here is not the first non-relativistic heterotic supergravity in the literature, it is very likely that all these approaches are connected by field redefinitions.
+In particular, in this work we also compute part of the finite Lagrangian of the formulation given in [21] for a particular case where the vielbein is gauge invariant (Appendix B).
+While there is not yet a proof, a possible way to connect these two formulations of non-relativistic heterotic supergravity is to trivialize the gauge transformation of the vielbein in both theories (without imposing extra constraints), which will need field redefinitions on the b-field and the gauge fields, leading to the same Green-Schwarz mechanism.
+A remarkable point of the construction here presented is that the Green-Schwarz mechanism for the b-field can be trivialized when one takes the NR limit and therefore it is possible to redefine the b-field to the new $\tilde{b}_{\mu\nu}$ as in (3.17), and its curvature $\bar{h}_{\mu\nu\rho}$, both of them are gauge invariant.
+While there is not yet proof, this cancellation of the Green-Schwarz transformation could indicate that the gauge anomaly could be trivially resolved under the NR limit.
+We will discuss this point in the outlook, as a promising future direction of work.
 5.1 Outlook
-As promising avenue for future work, it would be very interesting to explore the incorporation of $\alpha´$-corrections to both the bosonic and the heterotic supergravity  [^2] See [32] for a pedagogical review on this topic and references therein.
+This paper opens several promising directions for future work.
+First, it would be very interesting to study what happens to the equations of motion under the NR limit, especially from the DFT point of view.
+Even in the bosonic case, these equations are finite by construction (since the generalized metric is finite) and therefore there are no equations missing in the double geometry.
+However, using the strong constraint to break the duality group could be related to the analysis of the equations of motion performed in [13].
+When the analysis is performed directly from the NR supergravity perspective, the Poisson equation is missed and therefore it is not very clear how this equation is lost upon parametrization of the generalized metric.
+Once this mechanism is understood for the bosonic case, the inclusion of the gauge group should be straightforward.
+Second, a very promising line of research is behind the cancellation of the Green-Schwarz mechanism under the NR expansion shown in this paper.
+Using the field redefinition given in (3.17), we have shown that the Green-Schwarz mechanism for gauge transformations can be trivialized.
+The key point here is the presence of the vector $\alpha_{-i}$, instead of the connection $a_{\mu i}$, which allowed us to remove this transformation.
+This is a strong indication that the anomaly cancellation in the heterotic string could not need a fixed gauge group to work under the NR limit.
+If this is the case, the present results open the possibility of new families of heterotic strings, considering arbitrary gauge groups.
+A detailed analysis of the anomaly cancellation is therefore an important step towards the construction of NR heterotic string theory.
+Third, another potential continuation is to include ${\cal N}=1$ supersymmetry in this framework, extending the results in [15].
+The inclusion of this symmetry in heterotic DFT requires the use of the generalized vielbein and, therefore, it is not straightforward to construct a finite action in the double geometry.
+However, once this construction is understood, the inclusion of the gaugino dynamics will be obtained by promoting the duality group from $O(D,D)$ to $O(D,D+N)$.
+So, this is a worthy program to explore NR heterotic supergravity in the presence of fermions.
+Finally, a very ambitious continuation would be the incorporation of $\alpha$´-corrections to both the bosonic and the heterotic supergravity  [^2] See [32] for a pedagogical review on this topic and references therein.
 in the NR limit.
-Since the Riemann tensor contains divergences, one should study whether these are controlled by the $\hat{H}_{\mu\nu\rho}$ (or $\bar{H}_{\mu\nu\rho}$ in the heterotic case) contributions, as happens at the two derivative level.
+Since the Riemann tensor contains divergences, one should study whether these are controlled by the $\hat{H}_{\mu\nu\rho}$ (or $\bar{H}_{\mu\nu\rho}$ in the heterotic case) contributions, as happens at the two-derivative level.
 The covariant way of writing the Ricci scalar using the covariant derivative of $\tau_{\mu}{}^{a}$ and $\tau^{\mu}{}_{a}$ could be easily extended in order to incorporate the spin connection $\omega_{\mu ab}$, so that the action is also invariant under the $SO(1,1)$ transformations.
 In this way, the higher-derivative contributions could be written in a more compact form, since we expect large contributions if the curvatures are not chosen in an optimal way.
-While DFT has been proved a powerful tool to read the expansion of the A-field [22], taking the NR limit at the DFT level including higher-derivative terms is probably the best strategy to read these corrections and then parametrize them to the supergravity level.
+While DFT has been proved to be a powerful tool for reading the expansion of the A-field [22], taking the NR limit at the DFT level including higher-derivative terms is probably the best strategy to read these corrections and then parametrize them at the supergravity level.
+While in this paper we show that the Green-Schwarz mechanism for the gauge symmetry can be trivialized using field redefinitions, it is not clear that the same should happen for the gravitational one when addressing the $\alpha^{\prime}$-corrections for heterotic supergravity under the non-relativistic limit.
 Acknowledgements
-E.L. is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.
-Appendix A Details on the cancellation of divergences: gauge case
-The field strength $F_{\mu\nu}{}^{i}$, after expanding $A_{\mu}{}^{i}$ considering (1.12) is given by
-$F^{(1)}_{\mu\nu}{}^{i}=\alpha_{-}^{i}\partial_{\mu}{\tau_{\nu}^{-}}+\partial_{\mu}{\alpha_{-}^{i}}\tau_{\nu}^{-}-\alpha_{-}^{i}\partial_{\nu}{\tau_{\mu}^{-}}-\partial_{\nu}{\alpha_{-}^{i}}\tau_{\mu}^{-}\,,$ (A.1)
-$F^{(0)}_{\mu\nu}{}^{i}=+\alpha_{-}^{j}\eta^{ik}\tau_{\mu}^{-}a_{\nu}^{i1}f_{jki1}-\alpha_{-}^{j}\eta^{ik}\tau_{\nu}^{-}a_{\mu}^{i1}f_{jki1}\,,$ (A.2)
-$F^{(-1)}_{\mu\nu}{}^{i}=\partial_{\mu}{a_{\nu}^{i}}-\partial_{\nu}{a_{\mu}^{i}}\,,$ (A.3)
-$F^{(-2)}_{\mu\nu}{}^{i}=-\eta^{ij}a_{\mu}^{k}a_{\nu}^{i1}f_{jki1}\,,$ (A.4)
-where the notation $F^{(i)}_{\mu\nu}{}^{i}$ denotes the terms of order $c^{i}$ in the expansion.
-The contraction of this field strength with two inverse metrics, $F^{\mu\nu i}=F_{\rho\sigma}g^{\mu\rho}g^{\nu\sigma}$ yields the expression,
-$F^{\mu\nu i}=F^{(1)\mu\nu i}+\mathcal{O}\big{(}\frac{1}{c}\big{)}\,,$ (A.5)
+The author is supported by the SONATA BIS grant 2021/42/E/ST2/00304 from the National Science Centre (NCN), Poland.
+Appendix A Details on the cancellation of divergences when $\hat{A}_{\mu}$ is included
+The field strength $\hat{F}_{\mu\nu}{}^{i}$, after expanding $\hat{A}_{\mu}{}^{i}$ considering (1.15), is given by
+$\hat{F}^{(1)}_{\mu\nu}{}^{i}=\alpha_{-}^{i}\partial_{\mu}{\tau_{\nu}^{-}}+\partial_{\mu}{\alpha_{-}^{i}}\tau_{\nu}^{-}-\alpha_{-}^{i}\partial_{\nu}{\tau_{\mu}^{-}}-\partial_{\nu}{\alpha_{-}^{i}}\tau_{\mu}^{-}\,$, (A.1)
+$\hat{F}^{(0)}_{\mu\nu}{}^{i}=+\alpha_{-}^{j}\eta^{ik}\tau_{\mu}^{-}a_{\nu}^{l}f_{jkl}-\alpha_{-}^{j}\eta^{ik}\tau_{\nu}^{-}a_{\mu}^{l}f_{jkl}\,$, (A.2)
+$\hat{F}^{(-1)}_{\mu\nu}{}^{i}=\partial_{\mu}{a_{\nu}^{i}}-\partial_{\nu}{a_{\mu}^{i}}\,$, (A.3)
+$\hat{F}^{(-2)}_{\mu\nu}{}^{i}=-\eta^{ij}a_{\mu}^{k}a_{\nu}^{l}f_{jkl}\,$, (A.4)
+where the notation $\hat{F}^{(i)}_{\mu\nu}$ denotes the terms of order $c^{i}$ in the expansion.
+The contraction of this field strength with two inverse metrics, $\hat{F}^{\mu\nu i}=\hat{F}_{\rho\sigma}{}^{i}\hat{g}^{\mu\rho}\hat{g}^{\nu\sigma}$ yields the expression
+$\hat{F}^{\mu\nu i}=\hat{F}^{(1)\mu\nu i}+\mathcal{O}\big(\frac{1}{c}\big)\,$, (A.5)
 with
-$F^{(1)\mu\nu i}=\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\rho}h^{\nu\sigma}-\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\sigma}h^{\nu\rho}\,.$ (A.6)
-As the reader can easily check, the only contraction which produces non-vanishing divergences is given by $F_{(1)\mu\nu}{}^{i}F^{(1)\mu\nu j}\eta_{ij}$, which leads to the result (2.5).
-For the $-\frac{1}{12}\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho}$ contributions, an important observation is that
+$\hat{F}^{(1)\mu\nu i}=\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\rho}h^{\nu\sigma}-\alpha_{-}^{i}\partial_{\rho}{\tau_{\sigma}^{-}}h^{\mu\sigma}h^{\nu\rho}\,.$ (A.6)
+As the reader can easily check, the only contraction which produces non-vanishing divergences is given by $\hat{F}_{(1)\mu\nu}{}^{i}\hat{F}^{(1)\mu\nu j}\eta_{ij}$, which leads to the result (2.5).
+Before analyzing the $\bar{H}$ squared contribution, we recall the form of the Chern-Simons terms,
+$\bar{H}_{\mu\nu\rho}=3\left(\partial_{[\mu}\hat{B}_{\nu\rho]}-\hat{C}_{\mu\nu\rho}^{(g)}\right)\,$, (A.7)
+$\hat{C}_{\mu\nu\rho}^{(g)}=\hat{A}^{i}_{[\mu}\partial_{\nu}\hat{A}_{\rho]i}-\frac{1}{3}\hat{f}_{ijk}\hat{A}_{\mu}^{i}\hat{A}_{\nu}^{j}\hat{A}_{\rho}^{k}\,.$ (A.8)
+For the $\frac{1}{12}\bar{H}_{\mu\nu\rho}\bar{H}^{\mu\nu\rho}$ contributions, an important observation is that
 $\bar{H}_{\mu\nu\rho}=\bar{H}^{(2)}_{\mu\nu\rho}+\bar{H}_{(0)\mu\nu\rho}+\mathcal{O}\big(\frac{1}{c}\big)\,$, (A.9)
-while the same object when all the indices raised by the inverse metric has the following form,
+while the same object, when all the indices raised by the inverse metric, has the following form
 $\bar{H}^{\mu\nu\rho}=\bar{H}^{(0)\mu\nu\rho}+\mathcal{O}\big(\frac{1}{c^{2}}\big)\,.$ (A.10)
@@
 $+\frac{1}{2}\alpha_{-}^{i}\eta_{ij}\partial_{\sigma}{\tau_{\gamma}^{-}}a_{\epsilon}^{j}h^{\mu\gamma}h^{\nu\sigma}h^{\rho\epsilon}\,.$
-The multiplication of the previous two quantities leads directly to the result (2.6).
+The multiplication of the previous two quantities leads directly to the result (2.6) after imposing the Newton-Cartan relations (1.6).
+Appendix B The Bergshoeff-Romano action
+The finite action from the construction given in [21] in the case $v_{-}{}^{i}=0$ is given by
+$L=ee^{-2\varphi}\Big(L_{\textrm{u}ngauged}+T_{H^{2}}+T_{F^{2}}\Big)$ (B.1)
+where the gauged contributions of the Lagrangian are given by (we use A,B… for the transversal indices),
+$T_{H^{2}}=\frac{1}{2}\alpha_{+}^{i}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{b_{\sigma\gamma}}a_{A}^{j}e_{\epsilon}^{A}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\frac{1}{2}\eta_{ij}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{e_{\gamma}^{A}}a_{A}^{i}a_{B}^{j}e_{\epsilon}^{B}h^{\mu\epsilon}h^{\nu\sigma}h^{\rho\gamma}$
+$+\frac{1}{2}\eta_{ij}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{b_{\rho\sigma}}a_{B}^{j}e_{\gamma}^{A}e_{\epsilon}^{B}h^{\mu\rho}h^{\nu\epsilon}h^{\sigma\gamma}-\frac{1}{2}\alpha_{+}^{i}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{b_{\sigma\gamma}}a_{A}^{j}e_{\epsilon}^{A}h^{\mu\sigma}h^{\nu\rho}h^{\gamma\epsilon}$
+$-\frac{1}{2}\eta_{ij}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{e_{\gamma}^{A}}a_{A}^{i}a_{B}^{j}e_{\epsilon}^{B}h^{\mu\gamma}h^{\nu\sigma}h^{\rho\epsilon}-\frac{1}{2}\eta_{ij}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{b_{\rho\sigma}}a_{B}^{j}e_{\gamma}^{A}e_{\epsilon}^{B}h^{\mu\rho}h^{\nu\gamma}h^{\sigma\epsilon}$
+$+\frac{1}{2}\alpha_{+}^{i}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{b_{\sigma\gamma}}a_{A}^{j}e_{\epsilon}^{A}h^{\mu\rho}h^{\nu\sigma}h^{\gamma\epsilon}+\frac{1}{2}\eta_{ij}\partial_{\mu}{b_{\nu\rho}}\partial_{\sigma}{e_{\gamma}^{A}}a_{A}^{i}a_{B}^{j}e_{\epsilon}^{B}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$
+$+\frac{1}{2}\eta_{ij}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{b_{\rho\sigma}}a_{B}^{j}e_{\gamma}^{A}e_{\epsilon}^{B}h^{\mu\nu}h^{\rho\gamma}h^{\sigma\epsilon}-\frac{1}{2}\partial_{\mu}{b_{\nu\rho}}a_{A}^{i}a_{B}^{j}a_{C}^{k}e_{\sigma}^{A}e_{\gamma}^{B}e_{\epsilon}^{C}f_{ijk}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$
+$-\frac{1}{2}\alpha_{+}^{i}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{\tau_{\sigma}^{+}}\tau^{\sigma}_{c}a_{A}^{j}e_{\gamma}^{A}h^{\mu\gamma}h^{\nu\rho}-\frac{1}{2}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\alpha_{+}^{i}}\partial_{\nu}{\tau_{\rho}^{b}}\tau_{\sigma}^{+}\tau^{\sigma}_{c}a_{A}^{j}e_{\gamma}^{A}h^{\mu\rho}h^{\nu\gamma}$
+$-\frac{1}{2}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{e_{\sigma}^{A}}\tau^{\sigma}_{c}a_{A}^{i}a_{B}^{j}e_{\gamma}^{B}h^{\mu\gamma}h^{\nu\rho}-\frac{1}{2}\alpha_{+}^{i}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{\tau_{\sigma}^{+}}\tau^{\rho}_{c}a_{A}^{j}e_{\gamma}^{A}h^{\mu\sigma}h^{\nu\gamma}$
+$-\frac{1}{2}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{e_{\sigma}^{A}}\tau^{\rho}_{c}a_{A}^{i}a_{B}^{j}e_{\gamma}^{B}h^{\mu\sigma}h^{\nu\gamma}-\frac{1}{2}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{a_{A}^{i}}\tau^{\rho}_{c}a_{B}^{j}e_{\sigma}^{A}e_{\gamma}^{B}h^{\mu\sigma}h^{\nu\gamma}$
+$-\frac{1}{2}\alpha_{+}^{i}\alpha_{+}^{j}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{\tau_{\sigma}^{+}}\tau_{\gamma}^{+}\tau^{\gamma}_{c}h^{\mu\rho}h^{\nu\sigma}-\frac{1}{2}\alpha_{+}^{i}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{e_{\sigma}^{A}}\tau_{\gamma}^{+}\tau^{\gamma}_{c}a_{A}^{j}h^{\mu\rho}h^{\nu\sigma}$
+$-\frac{1}{2}\alpha_{+}^{i}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{a_{A}^{j}}\tau_{\sigma}^{+}\tau^{\sigma}_{c}e_{\gamma}^{A}h^{\mu\rho}h^{\nu\gamma}+\frac{1}{2}\alpha_{+}^{i}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{\tau_{\sigma}^{+}}\tau^{\rho}_{c}a_{A}^{j}e_{\gamma}^{A}h^{\mu\gamma}h^{\nu\sigma}$
+$+\frac{1}{2}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{e_{\sigma}^{A}}\tau^{\rho}_{c}a_{A}^{i}a_{B}^{j}e_{\gamma}^{B}h^{\mu\gamma}h^{\nu\sigma}+\frac{1}{2}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{a_{A}^{i}}\tau^{\rho}_{c}a_{B}^{j}e_{\sigma}^{A}e_{\gamma}^{B}h^{\mu\gamma}h^{\nu\sigma}$
+$+\frac{1}{2}\alpha_{+}^{i}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{\tau_{\sigma}^{+}}\tau^{\sigma}_{c}a_{A}^{j}e_{\gamma}^{A}h^{\mu\rho}h^{\nu\gamma}+\frac{1}{2}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\alpha_{+}^{i}}\partial_{\nu}{\tau_{\rho}^{b}}\tau_{\sigma}^{+}\tau^{\sigma}_{c}a_{A}^{j}e_{\gamma}^{A}h^{\mu\nu}h^{\rho\gamma}$
+$+\frac{1}{2}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{e_{\sigma}^{A}}\tau^{\sigma}_{c}a_{A}^{i}a_{B}^{j}e_{\gamma}^{B}h^{\mu\rho}h^{\nu\gamma}+\frac{1}{2}\alpha_{+}^{i}\alpha_{+}^{j}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{\tau_{\sigma}^{+}}\tau_{\gamma}^{+}\tau^{\gamma}_{c}h^{\mu\sigma}h^{\nu\rho}$
+$+\frac{1}{2}\alpha_{+}^{i}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{e_{\sigma}^{A}}\tau_{\gamma}^{+}\tau^{\gamma}_{c}a_{A}^{j}h^{\mu\sigma}h^{\nu\rho}+\frac{1}{2}\alpha_{+}^{i}\epsilon_{ab}\eta^{ac}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{b}}\partial_{\rho}{a_{A}^{j}}\tau_{\sigma}^{+}\tau^{\sigma}_{c}e_{\gamma}^{A}h^{\mu\gamma}h^{\nu\rho}$
+$+\alpha_{+}^{i}\epsilon_{ab}\eta^{ac}\partial_{\mu}{\tau_{\nu}^{b}}\tau_{\rho}^{+}\tau^{\rho}_{c}a_{A}^{j}a_{B}^{k}e_{\sigma}^{A}e_{\gamma}^{B}f_{ijk}h^{\mu\sigma}h^{\nu\gamma}-\frac{1}{8}\alpha_{+}^{i}\alpha_{+}^{j}\eta_{ik}\eta_{ji1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{\tau_{\sigma}^{+}}a_{A}^{k}a_{B}^{i1}e_{\gamma}^{A}e_{\epsilon}^{B}h^{\mu\rho}h^{\nu\sigma}h^{\gamma\epsilon}$
+$-\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{e_{\sigma}^{A}}a_{A}^{k}a_{B}^{j}a_{C}^{i1}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\rho}h^{\nu\sigma}h^{\gamma\epsilon}-\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{a_{A}^{k}}a_{B}^{j}a_{C}^{i1}e_{\sigma}^{A}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\rho}h^{\nu\sigma}h^{\gamma\epsilon}$
+$-\frac{1}{4}\alpha_{+}^{i}\alpha_{+}^{j}\eta_{ik}\eta_{ji1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{\tau_{\sigma}^{+}}a_{A}^{k}a_{B}^{i1}e_{\gamma}^{A}e_{\epsilon}^{B}h^{\mu\sigma}h^{\nu\epsilon}h^{\rho\gamma}-\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{e_{\sigma}^{A}}a_{A}^{k}a_{B}^{j}a_{C}^{i1}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\epsilon}h^{\nu\rho}h^{\sigma\gamma}$
+$-\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{a_{A}^{k}}a_{B}^{j}a_{C}^{i1}e_{\sigma}^{A}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\epsilon}h^{\nu\rho}h^{\sigma\gamma}-\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{e_{\sigma}^{A}}a_{A}^{k}a_{B}^{j}a_{C}^{i1}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\sigma}h^{\nu\epsilon}h^{\rho\gamma}$
+$-\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{a_{A}^{k}}a_{B}^{j}a_{C}^{i1}e_{\sigma}^{A}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\sigma}h^{\nu\epsilon}h^{\rho\gamma}+\frac{1}{8}\alpha_{+}^{i}\alpha_{+}^{j}\eta_{ik}\eta_{ji1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{\tau_{\sigma}^{+}}a_{A}^{k}a_{B}^{i1}e_{\gamma}^{A}e_{\epsilon}^{B}h^{\mu\sigma}h^{\nu\rho}h^{\gamma\epsilon}$
+$+\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{e_{\sigma}^{A}}a_{A}^{k}a_{B}^{j}a_{C}^{i1}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\sigma}h^{\nu\rho}h^{\gamma\epsilon}+\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{a_{A}^{k}}a_{B}^{j}a_{C}^{i1}e_{\sigma}^{A}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\sigma}h^{\nu\rho}h^{\gamma\epsilon}$
+$+\frac{1}{8}\alpha_{+}^{i}\alpha_{+}^{j}\eta_{ik}\eta_{ji1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{\tau_{\sigma}^{+}}a_{A}^{k}a_{B}^{i1}e_{\gamma}^{A}e_{\epsilon}^{B}h^{\mu\epsilon}h^{\nu\sigma}h^{\rho\gamma}+\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{e_{\sigma}^{A}}a_{A}^{k}a_{B}^{j}a_{C}^{i1}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\epsilon}h^{\nu\sigma}h^{\rho\gamma}$
+$+\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{a_{A}^{k}}a_{B}^{j}a_{C}^{i1}e_{\sigma}^{A}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\epsilon}h^{\nu\sigma}h^{\rho\gamma}+\frac{1}{8}\alpha_{+}^{i}\alpha_{+}^{j}\eta_{ik}\eta_{ji1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{\tau_{\sigma}^{+}}a_{A}^{k}a_{B}^{i1}e_{\gamma}^{A}e_{\epsilon}^{B}h^{\mu\rho}h^{\nu\epsilon}h^{\sigma\gamma}$
+$+\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{e_{\sigma}^{A}}a_{A}^{k}a_{B}^{j}a_{C}^{i1}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\rho}h^{\nu\epsilon}h^{\sigma\gamma}+\frac{1}{4}\alpha_{+}^{i}\eta_{ij}\eta_{ki1}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{a_{A}^{k}}a_{B}^{j}a_{C}^{i1}e_{\sigma}^{A}e_{\gamma}^{B}e_{\epsilon}^{C}h^{\mu\rho}h^{\nu\epsilon}h^{\sigma\gamma}$
+$+\frac{1}{2}\alpha_{+}^{i}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}a_{A}^{j}a_{B}^{k}a_{C}^{i1}a_{D}^{i2}e_{\rho}^{A}e_{\sigma}^{B}e_{\gamma}^{C}e_{\epsilon}^{D}f_{ki1i2}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{8}\eta_{ij}\eta_{ki1}\partial_{\mu}{e_{\nu}^{A}}\partial_{\rho}{e_{\sigma}^{B}}a_{A}^{i}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\rho}h^{\nu\sigma}h^{\gamma\epsilon}$
+$-\frac{1}{4}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{e_{\rho}^{B}}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\sigma}^{A}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\nu}h^{\rho\sigma}h^{\gamma\epsilon}-\frac{1}{4}\eta_{ij}\eta_{ki1}\partial_{\mu}{e_{\nu}^{A}}\partial_{\rho}{e_{\sigma}^{B}}a_{A}^{i}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\sigma}h^{\nu\epsilon}h^{\rho\gamma}$
+$-\frac{1}{4}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{e_{\rho}^{B}}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\sigma}^{A}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\rho}h^{\nu\gamma}h^{\sigma\epsilon}-\frac{1}{4}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{e_{\rho}^{B}}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\sigma}^{A}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\epsilon}h^{\nu\sigma}h^{\rho\gamma}$
+$+\frac{1}{8}\eta_{ij}\eta_{ki1}\partial_{\mu}{e_{\nu}^{A}}\partial_{\rho}{e_{\sigma}^{B}}a_{A}^{i}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\sigma}h^{\nu\rho}h^{\gamma\epsilon}+\frac{1}{4}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{e_{\rho}^{B}}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\sigma}^{A}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\rho}h^{\nu\sigma}h^{\gamma\epsilon}$
+$+\frac{1}{8}\eta_{ij}\eta_{ki1}\partial_{\mu}{e_{\nu}^{A}}\partial_{\rho}{e_{\sigma}^{B}}a_{A}^{i}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\epsilon}h^{\nu\sigma}h^{\rho\gamma}+\frac{1}{4}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{e_{\rho}^{B}}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\sigma}^{A}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\epsilon}h^{\nu\gamma}h^{\rho\sigma}$
+$+\frac{1}{8}\eta_{ij}\eta_{ki1}\partial_{\mu}{e_{\nu}^{A}}\partial_{\rho}{e_{\sigma}^{B}}a_{A}^{i}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\rho}h^{\nu\epsilon}h^{\sigma\gamma}+\frac{1}{4}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{e_{\rho}^{B}}a_{B}^{k}a_{C}^{j}a_{D}^{i1}e_{\sigma}^{A}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\nu}h^{\rho\gamma}h^{\sigma\epsilon}$
+$+\frac{1}{2}\eta_{ij}\partial_{\mu}{e_{\nu}^{A}}a_{A}^{i}a_{B}^{j}a_{C}^{k}a_{D}^{i1}a_{E}^{i2}e_{\rho}^{B}e_{\sigma}^{C}e_{\gamma}^{D}e_{\epsilon}^{E}f_{ki1i2}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}-\frac{1}{8}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{a_{B}^{k}}a_{C}^{j}a_{D}^{i1}e_{\rho}^{A}e_{\sigma}^{B}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\nu}h^{\rho\sigma}h^{\gamma\epsilon}$
+$-\frac{1}{4}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{a_{B}^{k}}a_{C}^{j}a_{D}^{i1}e_{\rho}^{A}e_{\sigma}^{B}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}+\frac{1}{8}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{a_{B}^{k}}a_{C}^{j}a_{D}^{i1}e_{\rho}^{A}e_{\sigma}^{B}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\sigma}h^{\nu\rho}h^{\gamma\epsilon}$
+$+\frac{1}{8}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{a_{B}^{k}}a_{C}^{j}a_{D}^{i1}e_{\rho}^{A}e_{\sigma}^{B}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\epsilon}h^{\nu\gamma}h^{\rho\sigma}+\frac{1}{8}\eta_{ij}\eta_{ki1}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{a_{B}^{k}}a_{C}^{j}a_{D}^{i1}e_{\rho}^{A}e_{\sigma}^{B}e_{\gamma}^{C}e_{\epsilon}^{D}h^{\mu\nu}h^{\rho\epsilon}h^{\sigma\gamma}$
+$+\frac{1}{2}\eta_{ij}\partial_{\mu}{a_{A}^{i}}a_{B}^{j}a_{C}^{k}a_{D}^{i1}a_{E}^{i2}e_{\nu}^{A}e_{\rho}^{B}e_{\sigma}^{C}e_{\gamma}^{D}e_{\epsilon}^{E}f_{ki1i2}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}$
+$-\frac{1}{12}a_{A}^{i}a_{B}^{j}a_{C}^{k}a_{D}^{i1}a_{E}^{i2}a_{F}^{i3}e_{\mu}^{A}e_{\nu}^{B}e_{\rho}^{C}e_{\sigma}^{D}e_{\gamma}^{E}e_{\epsilon}^{F}f_{ijk}f_{i1i2i3}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\epsilon}\,$, (B.2)
+and
+$T_{F^{2}}$ $=$ $-\frac{1}{2}\alpha_{+}^{i}\alpha_{+}^{j}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{\tau_{\sigma}^{+}}h^{\mu\rho}h^{\nu\sigma}-\alpha_{+}^{i}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{e_{\sigma}^{A}}a_{A}^{j}h^{\mu\rho}h^{\nu\sigma}-\alpha_{+}^{i}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{a_{A}^{j}}e_{\sigma}^{A}h^{\mu\rho}h^{\nu\sigma}$ (B.3)
+$+\frac{1}{2}\alpha_{+}^{i}\alpha_{+}^{j}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{\tau_{\sigma}^{+}}h^{\mu\sigma}h^{\nu\rho}+\alpha_{+}^{i}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{e_{\sigma}^{A}}a_{A}^{j}h^{\mu\sigma}h^{\nu\rho}+\alpha_{+}^{i}\eta_{ij}\partial_{\mu}{\tau_{\nu}^{+}}\partial_{\rho}{a_{A}^{j}}e_{\sigma}^{A}h^{\mu\sigma}h^{\nu\rho}$
+$+\alpha_{+}^{i}\partial_{\mu}{\tau_{\nu}^{+}}a_{A}^{j}a_{B}^{k}e_{\rho}^{A}e_{\sigma}^{B}f_{ijk}h^{\mu\rho}h^{\nu\sigma}-\frac{1}{2}\eta_{ij}\partial_{\mu}{e_{\nu}^{A}}\partial_{\rho}{e_{\sigma}^{B}}a_{A}^{i}a_{B}^{j}h^{\mu\rho}h^{\nu\sigma}-\eta_{ij}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{e_{\rho}^{B}}a_{B}^{j}e_{\sigma}^{A}h^{\mu\nu}h^{\rho\sigma}$
+$+\frac{1}{2}\eta_{ij}\partial_{\mu}{e_{\nu}^{A}}\partial_{\rho}{e_{\sigma}^{B}}a_{A}^{i}a_{B}^{j}h^{\mu\sigma}h^{\nu\rho}+\eta_{ij}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{e_{\rho}^{B}}a_{B}^{j}e_{\sigma}^{A}h^{\mu\rho}h^{\nu\sigma}+\partial_{\mu}{e_{\nu}^{A}}a_{A}^{i}a_{B}^{j}a_{C}^{k}e_{\rho}^{B}e_{\sigma}^{C}f_{ijk}h^{\mu\rho}h^{\nu\sigma}$
+$-\frac{1}{2}\eta_{ij}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{a_{B}^{j}}e_{\rho}^{A}e_{\sigma}^{B}h^{\mu\nu}h^{\rho\sigma}+\frac{1}{2}\eta_{ij}\partial_{\mu}{a_{A}^{i}}\partial_{\nu}{a_{B}^{j}}e_{\rho}^{A}e_{\sigma}^{B}h^{\mu\sigma}h^{\nu\rho}+\partial_{\mu}{a_{A}^{i}}a_{B}^{j}a_{C}^{k}e_{\nu}^{A}e_{\rho}^{B}e_{\sigma}^{C}f_{ijk}h^{\mu\rho}h^{\nu\sigma}$
+$-\frac{1}{4}\eta^{ij}a_{A}^{k}a_{B}^{i1}a_{C}^{i2}a_{D}^{i3}e_{\mu}^{A}e_{\nu}^{B}e_{\rho}^{C}e_{\sigma}^{D}f_{iki1}f_{ji2i3}h^{\mu\rho}h^{\nu\sigma}\,$,
+where $a_{A}^{i}=v_{A}^{i}$ and $\alpha_{+}^{i}=v_{+}^{i}$ transform both as gauge connections.
+Appendix C Demostration of covariance of $R^{(0)}$
+The difference between the variation of $R^{0}$ using the infinitesimal diffeomorphism transformations,
+$\delta_{\xi}\tau_{\mu}{}^{a}$ $=$ $\xi^{\rho}\partial_{\rho}{\tau_{\mu}{}^{a}}+\partial_{\mu}{\xi^{\rho}}\tau_{\rho}{}^{a}\,$,
+$\delta_{\xi}\tau^{\mu}{}_{a}$ $=$ $\xi^{\rho}\partial_{\rho}{\tau^{\mu}{}_{a}}-\partial_{\rho}{\xi^{\mu}}\tau^{\rho}{}_{a}\,$,
+$\delta_{\xi}h^{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}{h^{\mu\nu}}–2\partial_{\rho}{\xi^{(\mu}}h^{\nu)\rho}\,$,
+$\delta_{\xi}h_{\mu\nu}$ $=$ $\xi^{\rho}\partial_{\rho}{h_{\mu\nu}}+2\partial_{(\mu}{\xi^{\rho}}h_{\nu)\rho}\,$, (C.1)
+and $\xi^{\rho}\partial_{\rho}R^{(0)}$ (scalar transformation) gives
+$2\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\epsilon}}h_{\epsilon\delta}h^{\mu\nu}h^{\rho\delta}h^{\sigma\gamma}+2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau^{\mu}_{a}h_{\gamma\epsilon}h^{\nu\epsilon}h^{\rho\sigma}-2\partial_{\mu}{\tau^{\nu}_{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau_{\gamma}^{a}h_{\nu\epsilon}h^{\mu\epsilon}h^{\rho\sigma}$
+$+2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau_{\gamma}^{b}\tau^{\mu}_{a}\tau^{\nu}_{b}h^{\rho\sigma}-\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\epsilon}}h_{\epsilon\delta}h^{\mu\delta}h^{\nu\rho}h^{\sigma\gamma}-\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\epsilon}}\tau_{\epsilon}^{a}\tau^{\mu}_{a}h^{\nu\rho}h^{\sigma\gamma}$
+$-2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau^{\nu}_{a}h_{\gamma\epsilon}h^{\mu\epsilon}h^{\rho\sigma}-2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau_{\gamma}^{b}\tau^{\mu}_{b}\tau^{\nu}_{a}h^{\rho\sigma}-\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\mu}}h^{\nu\sigma}h^{\rho\gamma}$
+$-2\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\nu}}h^{\mu\rho}h^{\sigma\gamma}+2\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\nu}}h^{\mu\sigma}h^{\rho\gamma}-2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\mu}}\tau^{\rho}_{a}h^{\nu\sigma}+2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\nu}}\tau^{\rho}_{a}h^{\mu\sigma}$
+$-2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\nu}}\tau^{\mu}_{a}h^{\rho\sigma}+\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\mu}}h^{\nu\rho}h^{\sigma\gamma}+2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\mu}}\tau^{\nu}_{a}h^{\rho\sigma}-2\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\epsilon}}h_{\epsilon\delta}h^{\mu\sigma}h^{\nu\gamma}h^{\rho\delta}$
+$+2\partial_{\mu}{\tau^{\nu}_{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau_{\gamma}^{a}h_{\nu\epsilon}h^{\mu\rho}h^{\sigma\epsilon}-2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau^{\rho}_{a}h_{\gamma\epsilon}h^{\mu\sigma}h^{\nu\epsilon}-2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau_{\gamma}^{b}\tau^{\nu}_{b}\tau^{\rho}_{a}h^{\mu\sigma}$
+$+\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\epsilon}}h_{\epsilon\delta}h^{\mu\delta}h^{\nu\sigma}h^{\rho\gamma}+2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau^{\rho}_{a}h_{\gamma\epsilon}h^{\mu\epsilon}h^{\nu\sigma}+\partial_{\mu}{h_{\nu\rho}}\partial_{\sigma\gamma}{\xi^{\epsilon}}\tau_{\epsilon}^{a}\tau^{\mu}_{a}h^{\nu\sigma}h^{\rho\gamma}$
+$+2\partial_{\mu}{\tau_{\nu}^{a}}\partial_{\rho\sigma}{\xi^{\gamma}}\tau_{\gamma}^{b}\tau^{\mu}_{b}\tau^{\rho}_{a}h^{\nu\sigma}=0\,.$ (C.2)
+To obtain the last equality one needs to impose that $h_{\mu\nu}h^{\nu\rho}=\delta_{\mu}^{\rho}-\tau_{\mu}{}^{a}\tau^{\rho}{}_{a}$.